### v3

# Construction of 2-D Z-Complementary Array Code Sets with Flexible Lengths for Different System Requirements

Abstract
In this paper, we propose a new and optimal construction of two-dimensional (2-D) Z-complementary array code set (ZCACS) using multivariable extended Boolean functions (EBFs). The proposed 2-D arrays can be used in multi-carrier code division multiple access (MC-CDMA) systems. The proposed construction produces a better PMEPR upper bound than the existing constructions for such use cases. We also propose a tighter upper bound for the set size and showed that the proposed code set is optimal in special cases. Finally, We derive 2-D Golay complementary array set (GCAS) and Golay complementary set (GCS) from the proposed construction, which has application in uniform rectangular array (URA)-based massive multiple-input multiple-output (mMIMO) system to achieve omnidirectional transmission. The simulation result shows the performance benefits of the derived codes. In essence, we show that the flexibility of the parameters of the proposed 2-D ZCACS makes it a good candidate for a wide range of use cases in wireless communication.
Index Terms:
Z-complementary array code set (ZCACS), peak-to-mean envelope power ratio (PMEPR), multi-carrier code division multiple access (MC-CDMA), multiple-input multiple-output (MIMO), uniform rectangular array (URA), space-time block code (STBC), omnidirectional transmission.
I Introduction
The area of sequences design with good correlation properties for applications in wireless communication started with the discovery of the Golay complementary pairs (GCP), introduced in the seminal paper of Marcel J. E. Golay [1]. Leveraging the ideal auto-correlation sum spectrum of these sequences, researchers have found several applications of them in communication systems, for example, channel estimation [2], radar [3], reduction of peak-to-mean envelope power ratio (PMEPR) in orthogonal frequency division multiplexing (OFDM) [4, 5], etc. However, there are practical limitations on such sequences. For example, computer search shows that binary GCPs exist only for the lengths of the form $2^{\alpha}10^{\beta}26^{\gamma}$. To generalize GCP, Tseng et al. introduced the Golay complementary set (GCS) in 1972 [6], which has increased flexibility in sequence lengths. In the OFDM system, the length of the sequence determines the number of orthogonal sub-carriers. Nevertheless, this flexibility of lengths was achieved with the trade-off in an increased PMEPR, which is upper bounded by the number of sequences. A different direction of generalization was the introduction of zero-correlation zone (ZCZ) by Fan et al. [7], where the auto-correlation sum is zero in a specific zone of time-shifts. However, in a multiple-access system, an entire sequence set (also called code) is assigned to a user, and hence large numbers of codes are required which are orthogonal with respect to the correlation function. The research in this direction aims to achieve the optimality of the set size. For this purpose, mutually orthogonal Golay complementary set (MOGCS) [8], Z-complementary code set (ZCCS) [7], low-correlation zone (LCZ) sequence set [9], and quasi-complementary code set (QCCS)[10] were introduced.
Recently, two-dimensional (2-D) array sets and code sets with their 2-D correlation properties have attracted significant attention due to their practical application benefits, such as multi-carrier code division multiple access (MC-CDMA). The traditional one-dimensional (1-D) correlation functions are special cases of 2-D correlation functions by definition. Hence, it is possible to use 2-D arrays in place of 1-D sequences. In such cases, the primary issue is power control, which depends on the correlation property. The 2-D correlation property of 2-D array sets enables us to control PMEPR in one dimension and achieve interference-free transmission in the other dimension. Also, Turcsány et al. in [11, 12] showed that relaxing the orthogonality property of a 2-D MOGCS, 2-D code sets with larger set size can be used in 2D-MC-CDMA. The benefit of using 2-D arrays instead of 1-D sequences in such a system is the increased freedom to position the array elements in the time and frequency dimensions simultaneously.
Apart from it, some recent applications of 2-D array are the use of precoders for massive multiple-input multiple-output (mMIMO) to achieve omnidirectional transmission [13], ultra-wide band (UWB) communication [14, 15], etc. mMIMO has been extensively used in the 5th generation (5G) communication. It has the capability to increase the data rate by a large factor than the single-input single-output (SISO) system, making it an inseparable feature of future communication standards [16]. Hence, the literature has many works on 2-D code sets, such as 2-D Golay complementary array sets (GCAS) and 2-D quasi-complementary array sets (QCAS), for designing precoders in the mMIMO systems. The corresponding articles can be found here [17, 18, 19, 20, 21, 22]. However, applying 2-D QCAS as precoders has its limitations despite the increased flexibility of parameters. Because, the power radiation pattern is not omnidirectional here [22]. All this suggests that the existing 2-D codes does not completely solve the question of parameter flexibility in omnidirectional transmission. Another important parameter in such a system is the size of space-time block code (STBC) for encoding data [17]. Orthogonal STBCs with lower sizes are suitable for reducing complexity. It is an extremely difficult task to accommodate both MIMO and STBC configurations in a single construction of codes, due to the trade-off between the parameters. The best solution appears to be finding codes according to the requirements.
In this paper, as the main contribution, we propose 2-D Z-complementary array code sets (ZCACS) of array size $\left(2^{m}p\right)\times\left(2p_{1}^{m_{1}}p_{2}^{m_{2}}\dots p_{k}^{m_{k}}\right)$, where $m_{\alpha}\geq 2,\forall\alpha$, $p$, $p_{\alpha}$’s are prime numbers and $m\geq 1$. To do this, we first propose a construction of inter-group complementary (IGC) code set, which provides the necessary EBFs for the main construction. Then we analyze the PMEPR bounds and show that they are tighter than the existing constructions. We also investigate the set size upper bound and show that the proposed code set is optimal with respect to the bound for $p=2$. In addition to that, we derive suitable precoding matrices from the 2-D ZCACs for use in mMIMO to achieve omnidirectional transmission. Simulations are done to show the performance benefit of the precoders and parameter flexibility in URA and STBC.
The rest of the paper is organized as follows: in Section II, the preliminary definitions and lemmas are discussed. In Section III, the main construction is proposed. In Section IV, the bound for PMEPR is analyzed. In Section V, the set size bound of 2-D ZCACS is investigated. In Section VI, the application benefit of 2-D code sets in mMIMO is discussed in detail. In Section VII, we compare the proposed construction with previous works. Finally, in Section VIII, we draw conclusion.
II Preliminaries
In this section, we provide the definitions of the sequence/2-D array sets and code sets related to the work. Also, we give a structure of the EBFs used in the construction.
II-A Sequences and 2-D Arrays
The sequences and 2-D arrays considered here are assumed to be unimodular, polyphase and complex-valued, which are used in wireless systems with $q$-ary phase shift keying ($q$-PSK) modulation. A 2-D array $\mathbf{X}$ of size $L_{1}\times L_{2}$ is denoted by $X_{i,j}$, where $0\leq i<L_{1}$, $0\leq j<L_{2}$. For $L_{2}=1$, it becomes a sequence of length $L_{1}$, denoted by a lowercase bold letter.
Definition 1 (2-D ACCF )
The 2-D aperiodic cross-correlation function (ACCF) of two arrays $\mathbf{X}$ and $\mathbf{Y}$ of size $(L_{1},L_{2})$ at time shift $(\tau_{1},\tau_{2})$ is defined by
$C(\mathbf{X},\mathbf{Y})(\tau_{1},\tau_{2})=\begin{cases}&\sum\limits_{i=0}^{L_{1}-1}\sum\limits_{j=0}^{L_{2}-1}X_{i,j}Y^{*}_{i+\tau_{1},j+\tau_{2}},\\ \end{cases}$ (1)
where $X_{i,j},Y_{i,j}=0$ for $i\notin\{0,1,\dots,L_{1}-1\}$ or $j\notin\{0,1,\dots,L_{2}-1\}$. For $\mathbf{X}=\mathbf{Y}$, it is called 2-D aperiodic auto-correlation function (AACF), denoted by $A(\mathbf{X})(\tau_{1},\tau_{2})$.
Definition 2 (2-D Array Code)
An ordered set $\mathcal{X}=\{\mathbf{X}_{0},\mathbf{X}_{1},\dots,\mathbf{X}_{M-1}\}$ is called a 2-D array code of flock size $M$, where each $\mathbf{X}_{\eta}$ has size $L_{1}\times L_{2}$.
Definition 3 (2-D GCAS)
A set of 2-D arrays $\mathcal{X}=\{\mathbf{X}_{0},\mathbf{X}_{1},\dots,\mathbf{X}_{M-1}\}$ with size $L_{1}\times L_{2}$ is called a 2-D GCAS if they satisfy
$\sum_{\eta=0}^{M-1}A(\mathbf{X}_{\eta})(\tau_{1},\tau_{2})=\begin{cases}ML_{1}L_{2},&(\tau_{1},\tau_{2})=(0,0);\\ 0,&\text{otherwise}.\end{cases}$ (2)
For $M=2$, we call it a 2-D Golay complementary array pair (GCAP).
Definition 4 (2-D ZCAC ( [23]))
A set of 2-D arrays $\mathcal{X}=\{\mathbf{X}_{0},\mathbf{X}_{1},\dots,\mathbf{X}_{M-1}\}$ with size $L_{1}\times L_{2}$ is called a 2-D ZCAC if they satisfy
$\sum_{\eta=0}^{M-1}A(\mathbf{X}_{\eta})(\tau_{1},\tau_{2})=\begin{cases}ML_{1}L_{2},&(\tau_{1},\tau_{2})=(0,0);\\ 0,&0\leq\left|\tau_{1}\right|<Z_{1},\\ &0\leq\left|\tau_{2}\right|<Z_{2},\\ &(\tau_{1},\tau_{2})\neq(0,0).\end{cases}$ (3)
where $Z_{1}\leq L_{1}$, $Z_{2}\leq L_{2}$. Here, $Z_{1}\times Z_{2}$ is called the rectangular ZCZ size, $M$ is called the flock size.
Note 1
When $M=2$, it is called a 2-D Z-complementary array pair (ZCAP).
Definition 5 (2-D ZCACS ([23]))
A collection of 2-D ZCAC $\mathcal{Z}=\{\mathcal{X}_{0},\mathcal{X}_{1},\dots,\mathcal{X}_{K-1}\}$ where $\mathcal{X}_{\nu}=\{\mathbf{X}_{0}^{\nu},\mathbf{X}_{1}^{\nu},\dots,\mathbf{X}_{M-1}^{\nu}\}$ and each $\mathbf{X}_{i}^{\nu}$ has equal size of $L_{1}\times L_{2}$, is called a 2-D ZCACS if they satisfy
$\begin{split}&\sum_{\eta=0}^{M-1}C(\mathbf{X}_{\eta}^{\nu_{1}},\mathbf{X}_{\eta}^{\nu_{2}})(\tau_{1},\tau_{2})\\ &=\begin{cases}ML_{1}L_{2},&\nu_{1}=\nu_{2},(\tau_{1},\tau_{2})=(0,0);\\ 0,&\nu_{1}=\nu_{2},(\tau_{1},\tau_{2})\neq(0,0),\\ &0\leq\left|\tau_{1}\right|<Z_{1},0\leq\left|\tau_{2}\right|<Z_{2};\\ 0,&\nu_{1}\neq\nu_{2},\left|\tau_{1}\right|<Z_{1},\left|\tau_{2}\right|<Z_{2};\\ \end{cases}\end{split}$ (4)
[…]
The known theoretical upper bound on the set size of a 2-D ZCACS is provided below.
Lemma 1 ([23])
[…]
$K\leq\frac{M(L_{1}+Z_{1}-1)(L_{2}+Z_{2}-1)}{Z_{1}Z_{2}}.$ (5)
A 2-D ZCACS is called optimal if $K$ achieves its upper bound.
Remark 1
1-D correlation functions and sequence sets appear as a special case of this 2-D versions. The corresponding definitions follow analogously with the change in $L_{1}=L,L_{2}=1$ and $Z_{1}=Z,Z_{2}=1$. In 1-D case, 2-D GCAS, GCAP, ZCAC, ZCAP, ZCACS are called Golay complementary set (GCS), GCP, Z-complementary set (ZCS), Z-complementary pair (ZCP), and Z-complementary code set (ZCCS), respectively.
Lemma 2 ([24, 25])
[…]
$K\leq M\left(\frac{L+Z-1}{Z}\right).$ (6)
[…]
Now, we define some special sequence sets which have been used in the construction process of 2-D ZCACS. We define those sequence sets in association with the 2-D counterparts as discussed in Remark 1.
Definition 6 (Complementary Mate)
Two GCPs $(\mathbf{a}_{1},\mathbf{b}_{1})$ and $(\mathbf{a}_{2},\mathbf{b}_{2})$ are called complementary mate if they satisfy
$C(\mathbf{a}_{1},\mathbf{a}_{2})(\tau)+C(\mathbf{b}_{1},\mathbf{b}_{2})(\tau)=0,\forall\tau.$ (7)
Definition 7 (IGC code set ([26]))
[…]
$\begin{split}\sum_{\eta=0}^{M-1}C(\mathbf{a}_{\eta}^{\nu_{1}},\mathbf{a}_{\eta}^{\nu_{2}})(\tau)=\begin{cases}ML,&\nu_{1}=\nu_{2},\tau=0;\\ 0,&\nu_{1}=\nu_{2},0<\left|\tau\right|<Z;\\ 0,&\nu_{1}\neq\nu_{2},\mathcal{C}_{\nu_{1}},\mathcal{C}_{\nu_{2}}\in\mathcal{S}^{g},\\ &\left|\tau\right|<Z;\\ 0,&\text{otherwise}.\end{cases}\\ \end{split}$ (8)
II-B Extended Boolean Function
Let $\mathbb{Z}_{n}=\{[0],[1],\dots,[n-1]\}$ be the ring of integers modulo $n$, where []delimited-[][$\cdot$][ ⋅ ] denote the residue class. Then any well-defined function $g:\mathbb{Z}_{L}\rightarrow\mathbb{Z}_{q}$ produces a sequence of length $L$, namely, $\left(g([0]),g([1]),\dots,g([L-1])\right)$ and it is the fundamental idea behind the sequence generation method using EBFs. Note that, any length $L$ can be written as $L=p_{1}^{m_{1}}p_{2}^{m_{2}}\dots p_{k}^{m_{k}}$, where $p_{\alpha}$’s are distinct primes. Then the Chinese Remainder Theorem (CRT) says that $\exists$ a ring isomorphism $\sigma$ such that
$\sigma:\mathbb{Z}_{L}\cong\mathbb{Z}_{p_{1}^{m_{1}}}\times\mathbb{Z}_{p_{2}^{m_{2}}}\times\dots\times\mathbb{Z}_{p_{k}^{m_{k}}},$ (9)
given by $x\mapsto\left(x\mod p_{1}^{m_{1}},x\mod p_{2}^{m_{2}},\dots,x\mod p_{k}^{m_{k}}\right)$. Now, for the purpose of maintaining the well-defined property of EBFs, we assume $\mathbb{Z}_{n}$ as the set $\{0,1,\dots,n-1\}$ of numbers only, dropping the residue-class property. The corresponding functions rules are set accordingly to map the elements of $\mathbb{Z}_{L}$ to $\mathbb{Z}_{q}$, which we show shortly.
We can define a bijection $\varphi:\mathbb{Z}_{p_{\alpha}^{m_{\alpha}}}\rightarrow\mathbb{Z}_{p_{\alpha}}^{m_{\alpha}}$ given by $\varphi(x)=\vec{x}=(x_{1},x_{2},\dots,x_{m_{\alpha}})$, where $\vec{x}$ is the $p_{\alpha}$-ary representation of the integer $x=\sum_{i=1}^{m_{\alpha}}x_{i}p_{\alpha}^{i-1}$. Hence, we get a bijective map
$\left(\varphi\circ\sigma\right):\mathbb{Z}_{L}\rightarrow\mathbb{Z}_{p_{1}}^{m_{1}}\times\mathbb{Z}_{p_{2}}^{m_{2}}\times\dots\times\mathbb{Z}_{p_{k}}^{m_{k}}.$ (10)
Now, for the co-domain of EBFs, we first take the set $\mathbb{Z}_{q}$ for any $q$. Then an EBF $f:\mathbb{Z}_{L}\rightarrow\mathbb{Z}_{q}$ can be thought of as a function
$f^{\prime}:\mathbb{Z}_{p_{1}}^{m_{1}}\times\mathbb{Z}_{p_{2}}^{m_{2}}\times\dots\times\mathbb{Z}_{p_{k}}^{m_{k}}\rightarrow\mathbb{Z}_{q},$ (11)
where $f(x)=f^{\prime}\left((\varphi\circ\sigma)(x)\right)$ and $f^{\prime}$ can be written as a function of $m^{\prime}=\sum_{\alpha=1}^{k}m_{\alpha}$ variables $\{v_{1},v_{2},\dots,v_{m^{\prime}}\}$. To make the function well-defined, we adopt the function rule as
$\left(x_{1},x_{2},\dots,x_{m^{\prime}}\right)\mapsto f^{\prime}_{\mathbb{Z}}(x_{1},x_{2},\dots,x_{m^{\prime}})\mod q,$ (12)
where $f^{\prime}_{\mathbb{Z}}(x_{1},x_{2},\dots,x_{m^{\prime}})$ denotes the value of the polynomial expression of $f^{\prime}(x_{1},x_{2},\dots,x_{m^{\prime}})$ in $\mathbb{Z}$.
Remark 2
Note that, if we apply a permutation function $\phi$ on $\sigma$, then the resultant function
$\left(\varphi\circ(\phi\circ\sigma)\right):\mathbb{Z}_{L}\rightarrow\mathbb{Z}_{p_{1}}^{m_{1}}\times\mathbb{Z}_{p_{2}}^{m_{2}}\times\dots\times\mathbb{Z}_{p_{k}}^{m_{k}}$ (13)
still offers a valid representations of the numbers $0\leq i<p_{1}^{m_{1}}p_{2}^{m_{2}}\dots p_{k}^{m_{k}}$, with some permutation on the original representation $\left(\varphi\circ\sigma\right)(i)$. For sequence generation in this paper, we shall need such permuted representation of $i$. This definition of EBFs allows us to permute the components in the direct product, i.e., $\mathbb{Z}_{p_{\alpha}}$’s, so that we have a more general form
$f^{\prime}:\mathbb{Z}_{p^{\prime}_{1}}^{m^{\prime}_{1}}\times\mathbb{Z}_{p^{\prime}_{2}}^{m^{\prime}_{2}}\times\dots\times\mathbb{Z}_{p^{\prime}_{k}}^{m^{\prime}_{k}}\rightarrow\mathbb{Z}_{q},$ (14)
where $p^{\prime}_{\alpha}$’s may not be distinct. We shall use this general form.
For the construction of 2-D ZCACS, we need the co-domain $\mathbb{Z}_{q}$ to be extended, which will allow some multiplicative invertible elements to appear in the range. In this process, we show that different kinds of such extensions are possible.
  1. (1)
    When considering $\mathbb{Z}_{q}$ as a ring, we may take the set $D=\{x:x\in\mathbb{Z}_{q},x^{n}\neq 0,\forall n\in\mathbb{N}\}$ such that $D$ is closed with respect to multiplication, and no nilpotent element exists in $D$. Then, we may construct the ring of fractions $D^{-1}\mathbb{Z}_{q}=\{xd^{-1}:x\in\mathbb{Z}_{q},d\in D\}$ by using localization, so that $\mathbb{Z}_{q}\subseteq D^{-1}\mathbb{Z}_{q}$ as a subring. This extension inherently provides well-defined binary operations in $D^{-1}\mathbb{Z}_{q}$. Note that such a $D$ always exists in $\mathbb{Z}_{q}$.
  2. (2)
    Or, we may consider the set $\mathbb{Z}_{q}=\{0,1,\dots,q-1\}$ as a set of integers only. Now, we can define a quotient set $\mathbb{Q}_{q}=\left\{\frac{a}{b}\mid a,b\in\mathbb{Z}_{q},b\neq 0\right\}$, where the addition and multiplication rule are set in such a way that they are well-defined and it avoids the degenerate case, i.e., $\mathbb{Q}_{q}=\{0\}$. This can be done by defining
    $(x_{1},x_{2},\dots,x_{m^{\prime}})\mapsto f^{\prime}_{\mathbb{Q}}(x_{1},x_{2},\dots,x_{m^{\prime}})$ (15)
    where $f^{\prime}_{\mathbb{Q}}(x_{1},x_{2},\dots,x_{m^{\prime}}):=\frac{a_{f^{\prime}}\mod q}{b_{f^{\prime}}\mod q}$, $a_{f^{\prime}},b_{f^{\prime}}\in\mathbb{Z}$. If it occurs that $b_{f^{\prime}}\mod q=0(\mod q)$, then we may define $f^{\prime}_{\mathbb{Q}}(x_{1},x_{2},\dots,x_{m^{\prime}}):=\kappa$, for some $\kappa\in\mathbb{Q}_{q}$.
For the sake of simplicity, we denote both types of extensions as $\mathbb{Q}_{q}$ and take either extension. It does not matter in the actual construction process.
From (14), we derive the associated complex-valued sequence corresponding to $f^{\prime}$. Let $v_{p^{\prime}_{\alpha},1},v_{p^{\prime}_{\alpha},2},\dots,v_{p^{\prime}_{\alpha},m_{\alpha}}$ denote the $m_{\alpha}$ variables which take values from $\mathbb{Z}_{p^{\prime}_{\alpha}}$, $\forall\alpha$. The radix-representation of $i$ is defined as $\left(\vec{i}_{1},\vec{i}_{2},\dots,\vec{i}_{k}\right)$. We can associate an EBF $f$ with a $\mathbb{Q}_{q}$-valued sequence of length $p_{1}^{m_{1}}p_{2}^{m_{2}}\dots p_{k}^{m_{k}}$ using
$\textbf{f}=\bigg{(}f\left(\vec{i}_{1},\vec{i}_{2},\dots,\vec{i}_{k}\right):0\leq i<{p^{\prime}}_{1}^{m_{1}}{p^{\prime}}_{2}^{m_{2}}\cdots{p^{\prime}}_{k}^{m_{k}}\bigg{)},$ (16)
where $i$ is varied in the following manner: First $\vec{i}_{1}$ is varied from 00 to ${p^{\prime}}_{1}^{m_{1}}-1$, keeping others fixed at zero-vector. Then both $\vec{i}_{2}$ and $\vec{i}_{1}$ are varied, keeping others fixed. In summary, at $l$-th step $\{\vec{i}_{l},\dots,\vec{i}_{1}\}$ are varied, keeping the rest fixed, for $l=1,2,\dots,k$. This gives an enumeration of the integers $0\leq i<{p^{\prime}}_{1}^{m_{1}}{p^{\prime}}_{2}^{m_{2}}\dots p^{\prime}_{k}{}^{m_{k}}$ similar to the Boolean table.
One can also associate a complex-valued sequence corresponding to an EBF $f$, which is given by
$\psi(f)=\bigg{(}\omega_{q}^{f\left(\vec{i}_{1},\vec{i}_{2},\dots,\vec{i}_{k}\right)}:0\leq i<{p^{\prime}}_{1}^{m_{1}}{p^{\prime}}_{2}^{m_{2}}\cdots{p^{\prime}}_{k}^{m_{k}}\bigg{)},$ (17)
where $\omega_{q}=\exp\left(\frac{2\pi\sqrt{-1}}{q}\right)$.
Note 2
For $p_{\alpha}=2,\forall\alpha$, this definition reduces to the definition of generalized Boolean function (GBF) $f:\mathbb{Z}_{2}^{m}\rightarrow\mathbb{Z}_{q}$.
II-C 2-D Extended Boolean Function
We further introduce some notations to produce 2-D arrays by using multivariable EBFs. Let $F_{\vec{d}}:\mathbb{Z}_{p_{1}}^{m_{1}}\times\mathbb{Z}_{p_{2}}^{m_{2}}\times\dots\times\mathbb{Z}_{p_{k}}^{m_{k}}\rightarrow\mathbb{Q}_{q}$ be a set of EBFs for $\vec{d}=(d_{p^{\prime}_{1}},d_{p^{\prime}_{2}},\dots,d_{p^{\prime}_{w}})\in\prod_{\alpha=1}^{w}\mathbb{Z}_{p^{\prime}_{\alpha}}$, where $p^{\prime}_{\alpha}$ are some distinct primes and $\vec{d}$ is the radix-representation of number $d$ for $0\leq d<p^{\prime}_{1}p^{\prime}_{2}\cdots p^{\prime}_{k}$. Let $K=\{F_{\vec{d}}:\vec{d}\in\prod_{\alpha=1}^{w}\mathbb{Z}_{p^{\prime}_{\alpha}}\}$ be an ordered set, which can be expressed as a $\mathbb{Q}_{q}$-valued 2-D array of size $L_{1}\times L_{2}$ by writing the corresponding sequences, where $L_{1}=p^{\prime}_{1}p^{\prime}_{2}\dots p^{\prime}_{w}$ and $L_{2}=p_{1}^{m_{1}}p_{2}^{m_{2}}\dots p_{k}^{m_{k}}$. The $d$-th row of $K$ is determined by the EBF $F_{\vec{d}}$.
Similar to the 1-D case, we can also associate a complex-valued matrix $\Psi(K)$ with $K$, where $\left(\Psi(K)\right)_{i,j}=\omega_{q}^{K_{i,j}}$ for all $i,j$.
TABLE I: 2-D array corresponding to the functions $F_{\vec{d}}$
$K$= $\vec{d}=(d_{2},d_{3})$ $(v_{2,1},v_{3,1}$) (0,0) (1,0) (0,1) (1,1) (0,2) (1,2) (0,0) 0 0 0 0 0 0 (1,0) 0 2 0 2 0 2 (0,1) 0 0 3 3 2 2 (1,1) 0 2 3 1 2 0 (0,2) 0 0 2 2 0 0 (1,2) 0 2 2 0 0 2
Example 1
Let $F_{\vec{d}}:\mathbb{Z}_{2}\times\mathbb{Z}_{3}\rightarrow\mathbb{Q}_{4}$ be defined by $F_{\vec{d}}(v_{2,1},v_{3,1})=2v_{2,1}d_{2}+3v_{3,1}d_{3}$, where $\vec{d}=(d_{2},d_{3})$. Then the $6\times 6$ 2-D array $K$ related to $F_{\vec{d}}$ is generated in TABLE I.
II-D Peak-to-Mean Envelope Power Ratio
We consider MC-CDMA system, which combines the principles of both CDMA and OFDM. In an asynchronous system, column sequence of 1-D code is considered for power control [27]. For a 2-D code, a 2-D array of that code can also be used, using any one of the dimensions for power control. Given a $\mathbb{Q}_{q}$-valued sequence $\mathbf{a}=\{a_{0},a_{1},\dots,a_{L-1}\}$, the transmitted signal can be modeled as the real part of the complex envelope $S_{\mathbf{a}}(t)=\sum_{i=0}^{L-1}\omega_{q}^{a_{i}+qf_{i}t}$, where $f_{i}=f+i\Delta f$, $f$ is a constant, $\Delta f$ is an integer multiple of symbol rate and $0\leq\Delta ft\leq 1$. The term $\frac{\mid S_{\mathbf{a}}(t)\mid^{2}}{L}$ is called the instantaneous-to-average power ratio (IAPR). The PMEPR of the sequence $\mathbf{a}$ is given by
$\text{PMEPR}(\mathbf{a})=\sup\limits_{0\leq\Delta ft\leq 1}\frac{\mid S_{\mathbf{a}}(t)\mid^{2}}{L}.$ (18)
In an ideal case, PMEPR should be as low as possible.
II-E mMIMO-based Omnidirectional Beamforming
In this subsection, we introduce the concept of precoding of mMIMO with URA for omnidirectional signal transmission. As described in [13, 28], we consider a downlink transmission from a base station (BS) equipped with URA to single-antenna user equipment (UE). We consider a $P\times Q$ URA of $PQ$ antennas. Let, $A(\phi,\theta)$ be the steering matrix corresponding to the URA at the direction angle $(\phi,\theta)$, where $\phi\in[0,\pi/2]$ and $\theta\in[0,2\pi]$. Then $A(\phi,\theta)$ is given by
$\begin{split}[A(\phi,\theta)]_{i,j}=e^{-\iota\frac{2\pi}{\lambda}id_{y}\sin\phi\sin\theta-\iota\frac{2\pi}{\lambda}jd_{x}\sin\phi\cos\theta},\end{split}$ (19)
where $\iota=\sqrt{-1}$, $10\leq i\leq P-$ 1, $10\leq j\leq Q-$ 1, $\lambda$ is the wavelength of the carrier, $d_{x}$ and $d_{y}$ are the inter-element distance of URA along the vertical and horizontal directions, respectively. At first, the data bits are modulated into a $K\times N$-STBC $\mathbf{S}$, the $(n,t)$-th entry given by $s_{n}(t)$, where $n=0,1,\dots,K-1$ and $t=0,1,\dots,N-1$ [29]. It is then beamformed with the precoding matrices $\mathbf{W}_{n}$ of size $P\times Q$ and sent over the wireless channel. The received signal in a line-of-sight (LOS) without multi-paths is given by
$y(t)=\sum_{n=0}^{K-1}\left[vec(A(\phi,\theta))^{T}\cdot vec(\mathbf{W}_{n})\right]s_{n}(t)+w(t),$ (20)
for $10\leq t\leq N-$ 1, where $vec(\cdot)$ denotes the vector produced by vertically stacking the columns of the corresponding matrix, $(\cdot)^{T}$ denotes the transpose of the matrix ()($\cdot$)( ⋅ ), and $w(t)$ is the additive white Gaussian noise (AWGN).
The next two lemmas show the usefulness of 1-D and 2-D codes used in such system for omnidirectional power property.
Lemma 3 ([28])
Let the precoding matrices $\{\mathbf{W}_{0},\mathbf{W}_{1},\dots,\mathbf{W}_{K-1}\}$ of size $P\times Q$ constitute a 2-D GCAS. Then the received power $\sum_{n=0}^{K-1}\left|\left[vec(A(\phi,\theta))^{T}\cdot vec(\mathbf{W}_{n})\right]\right|^{2}$ is independent of the direction $(\phi,\theta)$.
Lemma 4 ([13])
Let $\{\mathbf{c}_{0},\mathbf{c}_{1},\dots,\mathbf{c}_{P-1}\}$ be a GCS of length $Q$, with $P=K$. Then the received power $\sum_{n=0}^{K-1}\left|\left[vec(A(\phi,\theta))^{T}\cdot vec(\mathbf{W}_{n})\right]\right|^{2}$ is independent of the direction $(\phi,\theta)$ if $vec(\mathbf{W}_{n})$ is chosen to be
$vec(\mathbf{W}_{n})=\left[\underbrace{\mathbf{0}\cdots\mathbf{0}}_{n-\text{times}}~{}~{}~{}\mathbf{c}_{n}\underbrace{\mathbf{0}\cdots\mathbf{0}}_{(P-n-1)-\text{times}}\right]^{T},$ (21)
where $\mathbf{0}$ denotes the zero matrix of size $1\times Q$.
III Construction of 2-D ZCACS
In this section, we propose the construction of 2-D ZCACS along with its building block which is 2-D ZCAC. However, for this purpose, we first state a lemma that produces IGC code sets of flexible lengths of the form $p_{1}^{m_{1}-1}p_{2}^{m_{2}-1}\dots p_{k}^{m_{k}-1}(m_{\alpha}\geq 1,\forall\alpha)$.
III-A IGC Code Set
We take the EBFs $f_{\alpha}:\mathbb{Z}_{p_{\alpha}}^{m_{\alpha}-1}\rightarrow\mathbb{Q}_{q}$, for $\alpha=1,2,\dots,k$, where each $f_{\alpha}$ is defined by
$\begin{split}f_{\alpha}&=\frac{q}{p_{\alpha}}\sum_{\beta=1}^{m_{\alpha}-2}v_{p_{\alpha},\pi_{\alpha}(\beta)}v_{p_{\alpha},\pi_{\alpha}(\beta+1)}+\sum_{\beta=1}^{m_{\alpha}-1}c_{\alpha,\beta}v_{p_{\alpha},\beta}+c_{\alpha},\end{split}$ (22)
where $c_{\alpha,\beta},c_{\alpha}\in\mathbb{Q}_{q}$, $m_{\alpha}\geq 2$ for all $\alpha$, $\pi_{\alpha}$ is a permutation of the set $\{1,2,\dots,m_{\alpha}-1\}$, and $p_{\alpha}\mid q$ for all $\alpha$. Let the function
$f:\mathbb{Z}_{p_{1}}^{m_{1}-1}\times\mathbb{Z}_{p_{2}}^{m_{2}-1}\times\dots\times\mathbb{Z}_{p_{k}}^{m_{k}-1}\rightarrow\mathbb{Q}_{q}$ (23)
be defined by $f=f_{1}+f_{2}+\dots+f_{k}$. Now, for $\bm{\gamma}=(\gamma_{1},\gamma_{2},\dots,\gamma_{k}),\mathbf{s}=(s_{1},s_{2},\dots,s_{k}),\mathbf{t}=(t_{1},t_{2},\dots,t_{k})\in\mathbb{Z}_{p_{1}}\times\mathbb{Z}_{p_{2}}\times\dots\times\mathbb{Z}_{p_{k}}$, we define the set of functions
$a_{\mathbf{s},\mathbf{t}}^{\bm{\gamma}}:\mathbb{Z}_{p_{1}}^{m_{1}-1}\times\dots\times\mathbb{Z}_{p_{k}}^{m_{k}-1}\times\mathbb{Z}_{p_{1}}\times\dots\times\mathbb{Z}_{p_{k}}\rightarrow\mathbb{Q}_{q},$ (24)
given by $a_{\mathbf{s},\mathbf{t}}^{\bm{\gamma}}=R_{\mathbf{t}}^{\bm{\gamma}}+T_{\mathbf{s}}$, where $R_{\mathbf{t}}^{\bm{\gamma}}:\mathbb{Z}_{p_{1}}^{m_{1}-1}\times\dots\times\mathbb{Z}_{p_{k}}^{m_{k}-1}\rightarrow\mathbb{Q}_{q}$ is defined by
$R_{\mathbf{t}}^{\bm{\gamma}}=f+\sum_{\alpha=1}^{k}\frac{q}{p_{\alpha}}v_{p_{\alpha},\pi_{\alpha}(1)}\gamma_{\alpha}+\sum_{\alpha=1}^{k}\frac{q}{p_{\alpha}}v_{p_{\alpha},\pi_{\alpha}(m_{\alpha}-1)}t_{\alpha},$ (25)
and $T_{\mathbf{s}}:\mathbb{Z}_{p_{1}}\times\dots\times\mathbb{Z}_{p_{k}}\rightarrow\mathbb{Q}_{q}$ is defined by
$T_{\mathbf{s}}=\sum_{\alpha=1}^{k}\frac{q}{p_{\alpha}}v^{\prime}_{p_{\alpha}}s_{\alpha}.$ (26)
Here, $v^{\prime}_{p_{\alpha}}$’s are the variables corresponding to the last $\mathbb{Z}_{p_{\alpha}}$’s. Next, we define the code $\mathcal{C}_{\mathbf{s},\mathbf{t}}$ which is given by
$\mathcal{C}_{\mathbf{s},\mathbf{t}}=\begin{bmatrix}\psi(a_{\mathbf{s},\mathbf{t}}^{\bm{\gamma}})\\ \end{bmatrix}_{\bm{\gamma}\in\mathbb{Z}_{p_{1}}\times\mathbb{Z}_{p_{2}}\times\dots\times\mathbb{Z}_{p_{k}}},$ (27)
which is a $M\times L$ matrix with $M=p_{1}p_{2}\dots p_{k}$ and $L=p_{1}^{m_{1}}p_{2}^{m_{2}}\dots p_{k}^{m_{k}}$. The next lemma provides the construction of the IGC code set.
Lemma 5
The code set $\mathcal{S}=\{\mathcal{C}_{\mathbf{s},\mathbf{t}}:\mathbf{s},\mathbf{t}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}\}$ forms a $(K,M,L,Z)$-IGC code set with $K=(p_{1}p_{2}\dots p_{k})^{2}$, $M=p_{1}p_{2}\dots p_{k}$, $L=p_{1}^{m_{1}}p_{2}^{m_{2}}\dots p_{k}^{m_{k}}$ and $Z=p_{1}^{m_{1}-1}p_{2}^{m_{2}-1}\dots p_{k}^{m_{k}-1}$. The $\mathbf{t}$-th IGC code group is expressed as $I^{\mathbf{t}}=\{\mathcal{C}_{\mathbf{s},\mathbf{t}}:\mathbf{s}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}\}$.
Proof:
Please refer to Appendix A. ∎
We illustrate Lemma 5 by using an example.
Example 2
Let $p_{1}=2$, $p_{2}=3$, $q=6$ and $\pi_{1}(1)=1$, $\pi_{2}(1)=1$. Let the functions $f_{1}:\mathbb{Z}_{2}\rightarrow\mathbb{Q}_{6}$ and $f_{2}:\mathbb{Z}_{3}\rightarrow\mathbb{Q}_{6}$ be defined by $f_{1}=3v_{2,1}$ and $f_{2}=4v_{3,1}$. Then the function $a_{\mathbf{s},\mathbf{t}}^{\bm{\gamma}}:\mathbb{Z}_{2}\times\mathbb{Z}_{3}\times\mathbb{Z}_{2}\times\mathbb{Z}_{3}\rightarrow\mathbb{Q}_{6}$ is defined by $a_{\mathbf{s},\mathbf{t}}^{\bm{\gamma}}=f_{1}+f_{2}+(3v_{2,1}\gamma_{1}+2v_{3,1}\gamma_{2})+(3v_{2,1}t_{1}+2v_{3,1}t_{2})+(3v^{\prime}_{2}s_{1}+2v^{\prime}_{3}s_{2})$ where $\mathbf{s}=(s_{1},s_{2})$, $\mathbf{t}=(t_{1},t_{2})$, $\bm{\gamma}=(\gamma_{1},\gamma_{2})$ $\in\mathbb{Z}_{2}\times\mathbb{Z}_{3}$. Then the code set corresponding to functions $a_{\mathbf{s},\mathbf{t}}^{\bm{\gamma}}$ is an IGC code set of length $36$.
III-B Main Construction
Now, we come to the main construction of 2-D ZCAC and 2-D ZCACS using EBFs obtained from the construction of the IGC code set. Before that, let us state two more lemmas.
Lemma 6 ([30])
Let $q$ be an even positive integer and $f\left(x_{1},x_{2},\dots,x_{m}\right):\mathbb{Z}_{2}^{m}\rightarrow\mathbb{Z}_{q}$ be a function given by
$f=\frac{q}{2}\sum_{\beta=1}^{m-1}x_{\pi(\beta)}x_{\pi(\beta+1)}+\sum_{\beta=1}^{m}g_{\beta}x_{\beta}+e,$ (28)
where $\pi$ is a permutation of the set $\{1,2,\dots,m\}$ and $g_{\beta}\in\mathbb{Z}_{q}$. Then the complex-valued sequences $\mathbf{a}_{1}$ and $\mathbf{b}_{1}$, corresponding to the functions $a_{1}=f$ and $b_{1}=f+\frac{q}{2}x_{\pi(1)}$, respectively, forms a GCP of length $2^{m}$ for any $e\in\mathbb{Z}_{q}$.
Lemma 7 ([31])
Let $f$, $a_{1}$ and $b_{1}$ be defined as in Lemma 6. Then the complex-valued sequences $\mathbf{a}_{2}$ and $\mathbf{b}_{2}$, corresponding to the functions $a_{2}=\bar{f}+\frac{q}{2}x_{\pi(1)}$ and $b_{2}=\bar{f}$, respectively, forms a GCP of length $2^{m}$, where
$\bar{f}=\frac{q}{2}\sum_{\beta=1}^{m-1}\bar{x}_{\pi(\beta)}\bar{x}_{\pi(\beta+1)}+\sum_{\beta=1}^{m}g_{\beta}\bar{x}_{\beta}+e,$ (29)
and $\bar{x}_{i}=1-x_{i}$, $\forall i$. Moreover, $(\mathbf{a}_{1},\mathbf{b}_{1})$ and $(\mathbf{a}_{2},\mathbf{b}_{2})$ are complementary mates.
For further development, we let $2\mid q$, $a_{\mathbf{s},\mathbf{t}}^{\bm{\gamma}}$ be the same functions used in Lemma 5 and $p$ be a prime such that $p\mid q$. For $\vec{d}^{\prime}=(d_{1},d_{2},\dots,d_{m},\tilde{d}_{p})\in\mathbb{Z}_{2}^{m}\times\mathbb{Z}_{p}$, $\mathbf{t}_{1}\neq\mathbf{t}_{2}$, and $\bm{\zeta}=(\mathbf{s}_{1},\mathbf{s}_{2},\mathbf{t}_{1},\mathbf{t}_{2})$, we define a set of functions
$F_{\vec{d}^{\prime}}^{\bm{\gamma},\bm{\zeta}}:\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}^{m_{\alpha}-1}\times\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}\times\mathbb{Z}_{2}\rightarrow\mathbb{Q}_{q},$ (30)
given by
$\begin{split}F_{\vec{d}^{\prime}}^{\bm{\gamma},\bm{\zeta}}&=\bar{v^{\prime\prime}}\left\{a_{\mathbf{s}_{1},\mathbf{t}_{1}}^{\bm{\gamma}}+a_{1}(\vec{d})W_{1}(\tilde{d}_{p})+a_{2}(\vec{d})W_{2}(\tilde{d}_{p})\right\}\\ &+v^{\prime\prime}\left\{a_{\mathbf{s}_{2},\mathbf{t}_{2}}^{\bm{\gamma}}+b_{1}(\vec{d})W_{1}(\tilde{d}_{p})+b_{2}(\vec{d})W_{2}(\tilde{d}_{p})\right\},\end{split}$ (31)
where
$\begin{split}W_{i}(\tilde{d}_{p})=&\sum_{j=0}^{p-1}\frac{\prod_{i=0,i\neq j}^{p-1}(i-\tilde{d}_{p})}{\prod_{i=0,i\neq j}^{p-1}(i-j)}\delta^{i}_{j},\\ \end{split}$ (32)
$\begin{split}\delta^{1}_{j}=\begin{cases}1,&2\mid j;\\ 0,&2\not|~{}j;\end{cases}&\qquad\delta^{2}_{j}=\begin{cases}0,&2\mid j;\\ 1,&2\not|~{}j.\end{cases}\end{split}$ (33)
Here $v^{\prime\prime}$ denotes the variable corresponding to the last $\mathbb{Z}_{2}$ in (30), $\bar{v}^{\prime\prime}=1-v^{\prime\prime}$. $a_{1}$ and $b_{1}$ are the functions from Lemma 6, whereas $a_{2}$ and $b_{2}$ are the functions from Lemma 7. Also, we assumed $\vec{d}=(d_{1},d_{2},\dots,d_{m})\in\mathbb{Z}_{2}^{m}$.
Note 3
In the function $F_{\vec{d}^{\prime}}^{\bm{\gamma},\bm{\zeta}}$, we could take the sub-function of the form (32) only because we have extended the co-domain of EBF beforehand from $\mathbb{Z}_{q}$ to $\mathbb{Q}_{q}$.
The only thing that now remains is to determine the values of the parameter $\bm{\zeta}$, which will produce the set size of 2-D ZCACS. Let us define the set $\Omega=\{(\mathbf{s}_{1},\mathbf{s}_{2},\mathbf{t}_{1},\mathbf{t}_{2}):\mathbf{t}_{1}\neq\mathbf{t}_{2}\}$. We take a proper subset $\Gamma\subset\Omega$ such that for any two $\bm{\zeta}_{1}=(\mathbf{s}_{1}^{1},\mathbf{s}_{2}^{1},\mathbf{t}_{1}^{1},\mathbf{t}_{2}^{1}),\bm{\zeta}_{2}=(\mathbf{s}_{1}^{2},\mathbf{s}_{2}^{2},\mathbf{t}_{1}^{2},\mathbf{t}_{2}^{2})\in\Gamma$, $\mathbf{s}_{i}^{j},\mathbf{t}_{m}^{n}$ follow the condition for $1\leq i,j\leq 2$: (1) Either, $\mathbf{t}_{i}^{1}\neq\mathbf{t}_{j}^{2}$, (2) Or, $\mathbf{s}_{i}^{1}\neq\mathbf{s}_{j}^{2}$.
Following this, we consider only those functions $\mathcal{F}_{\vec{d}^{\prime}}^{\bm{\gamma},\bm{\zeta}}$ for which $\bm{\zeta}\in\Gamma$. Let the $\mathbb{Q}_{q}$-valued 2-D arrays be given by $\mathcal{K}^{\bm{\gamma},\bm{\zeta}}=\{F_{\vec{d}^{\prime}}^{\bm{\gamma},\bm{\zeta}}:\vec{d}^{\prime}\in\mathbb{Z}_{2}^{m}\times\mathbb{Z}_{p}\}$ for $\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}$ and we define $\mathbf{X}^{\bm{\gamma},\bm{\zeta}}=\Psi(\mathcal{K}^{\bm{\gamma},\bm{\zeta}})$.
Considering all these assumptions, we now propose the construction of the 2-D ZCAC and 2-D ZCACS.
Theorem 1
Each set $\mathcal{X}^{\bm{\zeta}}=\{\mathbf{X}^{\bm{\gamma},\bm{\zeta}}:\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}\}$ is a 2-D ZCAC of size $L_{1}\times L_{2}$, where $L_{1}=2^{m}p$ and $L_{2}=2p_{1}^{m_{1}}p_{2}^{m_{2}}\dots p_{k}^{m_{k}}$, ZCZ size is $Z_{1}\times Z_{2}$, where $Z_{1}=2^{m+1}$ and $Z_{2}=p_{1}^{m_{1}-1}p_{2}^{m_{2}-1}\dots p_{k}^{m_{k}-1}$ and flock size $M=p_{1}p_{2}\dots p_{k}$. The set $\{\mathcal{X}^{\bm{\zeta}}:\bm{\zeta}\in\Gamma\}$ is a 2-D ZCACS with set size $K=2p_{1}^{2}p_{2}^{2}\dots p_{k}^{2}$.
Proof:
Please refer to Appendix B. ∎
The following example illustrates Theorem 1.
Example 3
Let $p_{1}=2$, $p_{2}=3$, $q=6$, $\pi_{1}(1)=1$, $\pi_{2}(1)=1$, $a_{\mathbf{s}_{i}^{k},\mathbf{t}_{j}^{k}}^{\bm{\gamma}}$, $f_{1}$, and $f_{2}$ be the same functions as in Example 2. Let the function $f:\mathbb{Z}_{2}^{2}\rightarrow\mathbb{Z}_{6}$ be defined by $f=3x_{1}x_{2}+3$, where $\pi(1)=1,\pi(2)=2$. Then, let the functions $a_{1},b_{1},a_{2},b_{2}:\mathbb{Z}_{2}^{2}\rightarrow\mathbb{Z}_{6}$ be given by $a_{1}=3x_{1}x_{2}+3$, $b_{1}=3x_{1}x_{2}+3+3x_{1}$, and $a_{2},b_{2}$ are generated from Lemma 7. We also take $\bm{\zeta}_{1}=(\mathbf{s}_{1}^{1},\mathbf{s}_{2}^{1},\mathbf{t}_{1}^{1},\mathbf{t}_{2}^{1})$, where $\mathbf{s}_{1}^{1}=\mathbf{s}_{2}^{1}=(1,1)$ and $\mathbf{t}_{1}^{1}=(0,1)$, $\mathbf{t}_{2}^{1}=(1,2)$ and similarly we take $\bm{\zeta}_{2}=(\mathbf{s}_{1}^{2},\mathbf{s}_{2}^{2},\mathbf{t}_{1}^{2},\mathbf{t}_{2}^{2})$, where $\mathbf{s}_{1}^{2}=\mathbf{s}_{2}^{2}=(1,1)$ and $\mathbf{t}_{1}^{2}=(1,1)$, $\mathbf{t}_{2}^{2}=(1,0)$, so that $\mathbf{t}_{1}^{1},\mathbf{t}_{2}^{1},\mathbf{t}_{1}^{2}$ and $\mathbf{t}_{2}^{2}$ are all distinct. We also take $\vec{d}^{\prime}=(d_{1},d_{2},\tilde{d}_{p})\in\mathbb{Z}_{2}^{2}\times\mathbb{Z}_{3}$, $W_{1},W_{2}$ are generated from (32). Then, we get a 2-D ZCACS of size $12\times 72$ with set size $K=6$. The 2-D AACF sum of 2-D ZCAC $\{\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{1}}:\bm{\gamma}\in\mathbb{Z}_{2}\times\mathbb{Z}_{3}\}$ is given in the Fig. 1a. The size of the rectangular ZCZ is $8\times 6$. The plot of the 2-D ACCF sum of the set $\{\mathcal{X}^{\bm{\gamma},\bm{\zeta}_{1}},\mathcal{X}^{\bm{\gamma},\bm{\zeta}_{2}}\}$ is given in Fig. 1b.
Remark 3
It should be noted that, fixing $k=1$, and $p_{1}=2$, the set $\mathcal{X}=\{\mathbf{X}^{\gamma}:\gamma\in\mathbb{Z}_{2}\}$ becomes a 2-D ZCAP $\{\mathbf{X}^{0},\mathbf{X}^{1}\}$ of size $2^{m}p\times 2^{m_{1}+1}$ and rectangular ZCZ width is $2^{m+1}\times 2^{m_{1}-1}$. 2-D ZCAP [32, 33] of this size is reported for the first time in this paper, as a special case.
(a) The 2-D AACF sum of $\{\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{1}}:\bm{\gamma}\in\mathbb{Z}_{2}\times\mathbb{Z}_{3}\}$
(b) The 2-D ACCF sum of the set $\{\mathcal{X}^{\bm{\gamma},\bm{\zeta}_{1}},\mathcal{X}^{\bm{\gamma},\bm{\zeta}_{2}}\}$
Figure 1: Plot of 2-D ACCF/AACF sum corresponding to Example 3
The next two corollaries show that we can derive 2-D GCAS and GCS from the proposed construction of 2-D ZCAC, which can be used in mMIMO system.
Corollary 1
Let $\mathcal{X}^{\bm{\zeta}}$ be any 2-D ZCAC with $p=2$. Let $\bar{\mathcal{X}}^{\bm{\zeta}}$ be the set of 2-D arrays derived from $\mathcal{X}^{\bm{\zeta}}$, where each 2-D array is produced by deleting the following columns from $\mathcal{X}^{\bm{\zeta}}$: (i) From $(\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}+1)$-th to $(\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}})$-th column; (ii) From $(\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}+\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}+1)$-th to $(2\prod_{i=1}^{k}p_{\alpha}^{m_{\alpha}})$-th column. Then $\bar{\mathcal{X}}^{\bm{\zeta}}$ is a 2-D GCAS of size $(2^{m+1})\times(2p_{1}^{m_{1}-1}p_{2}^{m_{2}-1}\dots p_{k}^{m_{k}-1})$ and flock size $p_{1}p_{2}\dots p_{k}$.
Proof:
To prove this corollary, we use Lemma 5. Note that, after the truncation, the corresponding generating functions for the 2-D arrays of $\bar{\mathcal{X}}^{\bm{\zeta}}$ are given by
$\begin{split}F_{\vec{d}^{\prime}}^{\bm{\gamma},\bm{\zeta}}&=\bar{v^{\prime\prime}}\left\{a_{\mathbf{s}_{1},\mathbf{t}_{1}}^{\bm{\gamma}}+a_{1}(\vec{d})W_{1}(\tilde{d}_{p})+a_{2}(\vec{d})W_{2}(\tilde{d}_{p})\right\}\\ &+v^{\prime\prime}\left\{a_{\mathbf{s}_{2},\mathbf{t}_{2}}^{\bm{\gamma}}+b_{1}(\vec{d})W_{1}(\tilde{d}_{p})+b_{2}(\vec{d})W_{2}(\tilde{d}_{p})\right\},\end{split}$ (34)
where $a_{\mathbf{s}_{i},\mathbf{t}_{i}}^{\bm{\gamma}}=R_{\mathbf{t}_{i}}^{\bm{\gamma}}$, i.e., $T_{\mathbf{s}_{i}}=0$ for $i=1,2$. Now, proceeding similarly to the proof of Theorem 1, we get the result.
Corollary 2
Let $\mathbf{X}$ be any set of a 2-D ZCAC $\mathcal{X}^{\bm{\zeta}}$ with $p=2$. Then the column sequences $\{\mathbf{c}_{0},\mathbf{c}_{1},\dots,\mathbf{c}_{L_{1}-1}\}$ of $\mathbf{X}$ become a GCS of length $2^{m+1}$ and flock size $2p_{1}^{m_{1}}p_{2}^{m_{2}}\dots p_{k}^{m_{k}}$.
Proof:
This is a direct consequence of the construction of 2-D ZCAC proved in Theorem 1, by setting $p=2$. ∎
IV Analysis of PMEPR Bound
In this section, we investigate the upper bound of PMEPR for row and column sequences of the proposed 2-D ZCAC.
IV-A Row Sequence PMEPR
It is easy to observe that for the $m$-th row sequence $X_{m}^{row}$ of the 2-D array $\mathbf{X}$, we have
$\begin{split}\left|S_{X_{m}^{row}}(t)\right|^{2}=&L_{2}+\sum_{\tau_{2}\neq 0}A(X_{m}^{row})(\tau_{2})\omega_{q}^{-q\tau_{2}\Delta ft}.\end{split}$ (35)
For a fixed $\bm{\gamma}$ we take the $\mathbb{Q}_{q}$-valued 2-D array $K^{\bm{\gamma}}=\{F_{\tilde{\mathbf{d}}}^{\bm{\gamma},\zeta}:\tilde{\mathbf{d}}\in\mathbb{Z}_{2}^{m}\times\mathbb{Z}_{p}\}$. Now, from Lemma 9 used in the proof of Lemma 5, it can be deduced that
$\sum_{\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}A(K^{\bm{\gamma}}_{m})(\tau_{2})=0,$ (36)
for $\beta\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}<\left|\tau_{2}\right|<(\beta+1)\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1},\beta=0,1,\dots,2\prod_{\alpha=1}^{k}p_{\alpha}-1$ and $\left|\tau_{2}\right|=\delta\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1},\delta=\prod_{\alpha=1}^{k}p_{\alpha},\prod_{\alpha=1}^{k}p_{\alpha}+1,\dots,2\prod_{\alpha=1}^{k}p_{\alpha}-1$. From the above observation and the definition of AACF, we get
$\begin{split}&\left|\sum_{\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}A(K^{\bm{\gamma}}_{m})(\tau_{2})\right|\leq 2\left(\prod_{\alpha=1}^{k}p_{\alpha}-\beta\right)\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}},\end{split}$ (37)
for $\left|\tau_{2}\right|=\beta\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}$, where $\beta=1,2,\dots,\prod_{\alpha=1}^{k}p_{\alpha}-1$. Therefore, from (35), we can derive
$\begin{split}&\sum_{\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}\left|S_{K^{\bm{\gamma}}_{m}}(t)\right|^{2}\\ &\leq\!\!\!\!\!\!\!\sum_{\beta=1}^{\prod_{\alpha=1}^{k}p_{\alpha}-1}\!\!\!\!\!\!\sum_{\left|\tau_{2}\right|=\beta\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}}\left|\sum_{\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}\!\!\!\!\!\!\!\!\!A(K^{\bm{\gamma}}_{m})(\tau_{2})\right|+L_{2}\prod_{\alpha=1}^{k}p_{\alpha}\\ &=2\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}+1}\left(\prod_{\alpha}^{k}p_{\alpha}-1\right)+L_{2}\prod_{\alpha=1}^{k}p_{\alpha}.\end{split}$ (38)
But $\left|S_{K^{\bm{\gamma}}_{m}}(t)\right|^{2}\geq 0,\forall\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}$ and $L_{2}=2\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}$. So, we have
$\frac{\left|S_{K^{\bm{\gamma}}_{m}}(t)\right|^{2}}{L_{2}}\leq\prod_{\alpha}^{k}p_{\alpha}\left(\prod_{\alpha}^{k}p_{\alpha}-1\right)+\prod_{\alpha=1}^{k}p_{\alpha}.$ (39)
Hence, $\text{PMEPR}(K^{\bm{\gamma}}_{m})\leq\prod_{\alpha}^{k}p_{\alpha}\left(\prod_{\alpha}^{k}p_{\alpha}-1\right)+\prod_{\alpha=1}^{k}p_{\alpha}$. Similarly, we can compute the PMEPR upper bound of the column sequence.
IV-B Column Sequence PMEPR
Note that, similar to the previous case, for the $n$-th column sequence $X_{n}^{col}$, we have
$\begin{split}\left|S_{X_{n}^{col}}(t)\right|^{2}=&L_{1}+\sum_{\tau_{1}\neq 0}A(X_{n}^{col})(\tau_{1})\omega_{q}^{-q\tau_{1}\Delta ft}.\end{split}$ (40)
For this, we note that either ${K^{\bm{\gamma}}}^{col}_{n}=\mathcal{I}_{a}$ or ${K^{\bm{\gamma}}}^{col}_{n}=\mathcal{I}_{b}$, where $\mathcal{I}_{a}=a_{\mathbf{s}_{1},\mathbf{t}_{1}}^{\bm{\gamma}}+a_{1}(\vec{d})W_{1}(\tilde{d}_{p})+a_{2}(\vec{d})W_{2}(\tilde{d}_{p})$, $\mathcal{I}_{b}=a_{\mathbf{s}_{2},\mathbf{t}_{2}}^{\bm{\gamma}}+b_{1}(\vec{d})W_{1}(\tilde{d}_{p})+b_{2}(\vec{d})W_{2}(\tilde{d}_{p})$ for $(\vec{d},\tilde{d}_{p})\in\mathbb{Z}_{2}^{m}\times\mathbb{Z}_{p}$ and $a_{1},a_{2},b_{1},b_{2},W_{1}$ and $W_{2}$ are the functions used in Theorem 1. Note that for $p\neq 2$, we have
$\begin{split}&A(\mathcal{I}_{a})(\tau_{1})+A(\mathcal{I}_{b})(\tau_{1})\\ &=\begin{cases}(p-\delta)2^{m+1},&\left|\tau_{1}\right|=\delta 2^{m},\delta=0,2,\dots,p-1,\\ 0,&\text{otherwise}.\end{cases}\end{split}$ (41)
because $(a_{2},b_{2})$ is a complementary mate of $(a_{1},b_{1})$ by Lemma 7, and hence
$\begin{split}a_{1}(\vec{d})W_{1}(\tilde{d}_{p})+a_{2}(\vec{d})W_{2}(\tilde{d}_{p})=&\underbrace{a_{1}\otimes a_{2}\otimes\dots\otimes a_{1}}_{p\text{-times}},\\ b_{1}(\vec{d})W_{1}(\tilde{d}_{p})+b_{2}(\vec{d})W_{2}(\tilde{d}_{p})=&\underbrace{b_{1}\otimes b_{2}\otimes\dots\otimes b_{1}}_{p\text{-times}}.\end{split}$ (42)
For $p=2$, we have $a_{1}(\vec{d})W_{1}(\tilde{d}_{p})+a_{2}(\vec{d})W_{2}(\tilde{d}_{p})=a_{1}\otimes a_{2}$ and $b_{1}(\vec{d})W_{1}(\tilde{d}_{p})+b_{2}(\vec{d})W_{2}(\tilde{d}_{p})=b_{1}\otimes b_{2}$. In that case, $A(\mathcal{I}_{a})(\tau_{1})+A(\mathcal{I}_{b})(\tau_{1})=0,\forall\tau\neq 0$. So, arguing similarly to the previous case, we have
$\text{PMEPR}({K^{\bm{\gamma}}}^{col}_{n})\leq\begin{cases}2,&p=2;\\ p+\frac{1}{p},&p\neq 2.\end{cases}$ (43)
Remark 4
Take $p=2$, $k=1$ and $p_{1}=2$. Then Theorem 1 provides 2-D ZCAC of array size of the form $2^{n}\times 2^{m}$ and flock size $M=2$, which is comparable to the array size of 2-D GCAPs and 2-D GCASs in [34, Th. 12]. The PMEPR of the row and column sequences of 2-D GCAP and 2-D GCAS in [34] are upper bounded by $2^{v}$ for some positive integer $v\geq 1$. However, in our construction, for this case, the row and column sequence PMEPRs are upper bounded by $4$ and $2$, respectively. This shows that the proposed 2-D ZCAC emerges as a good candidate to reduce PMEPR compared to [34].
Now, we verify the claim made in Remark 4 by an explicit example.
Example 4
We take a 2-D ZCAC from Theorem 1 with the function of (31), where $a_{\mathbf{s_{i}},\mathbf{t_{j}}}^{\gamma}=v_{1}v_{2}+\gamma v_{1}+t_{j}v_{2}+s_{i}v_{3}$, $a_{1}=x_{1}x_{2}+1$, and $a_{2},b_{1},b_{2}$ are derived using Lemma 6 and Lemma 7. Then we get a 2-D ZCAC of size $8\times 16$ and flock size $2$. The row and column sequence PMEPRs are computed to be $3.4271$ and $2$, respectively. In Fig. 2, we have shown the corresponding plot of row and column sequence IAPR and PMEPR bound. We also have taken a binary $8\times 16$-GCAP from [34] by using the function $f=(z_{1}z_{2}+z_{2}z_{4}+z_{4}z_{5}+z_{5}z_{3}+z_{3}z_{6}+z_{6}z_{7})+1$ and plotted the corresponding PMEPR/IAPR curve in Fig. 2. It is evident from the figure that column sequence PMEPR of [34] is greater than $2$, for this particular example, while the proposed one is always bounded by $2$.
Figure 2: Row and column sequence IAPR/PMEPR
V Bound on the Set Size
[…]
Lemma 8 ([36])
Let $\mathcal{C}=\{\mathbf{A}^{\nu}:\nu=0,1,\dots,K-1\}$ be a set of $L_{1}\times L_{2}$ 2-D arrays with non-zero complex numbers elements of the form $\exp\left(\frac{2\pi n\sqrt{-1}}{q}\right)$. Let
$c_{max}=\max\limits_{\nu_{1}\neq\nu_{2}}\left|\sum_{i=0}^{L_{1}-1}\sum_{j=0}^{L_{2}-1}\mathbf{A}^{\nu_{1}}_{i,j}{\mathbf{A}^{\nu_{2}}}^{*}_{i,j}\right|.$ (44)
Then, we have
$\begin{split}K(K\!\!-1)c^{2}_{max}+\!K\left(L_{1}L_{2}\right)^{2}\!\geq\!\!\sum_{s_{i}=0}^{L_{1}-1}\sum_{s_{j}=0}^{L_{2}-1}\!\!\left(\sum_{\nu=0}^{K-1}\left|\!\mathbf{A}^{\nu}_{s_{i},s_{j}}\right|^{2}\right)^{2}.\end{split}$ (45)
With this lemma, we can state the tighter set size bound for 2-D ZCACS in the following theorem.
Theorem 2
[…]
$K\leq\left\lfloor\frac{ML_{1}L_{2}}{Z_{1}Z_{2}}\right\rfloor.$ (46)
Proof:
Please refer to Appendix C. ∎
Note 4
[…]
Remark 5
It is straightforward to verify that the bound proposed in Theorem 2 is tighter than the one derived in [23]. In our proposed construction, $L_{1}=2^{m}p$, $L_{2}=2\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}$, $Z_{1}=2^{m+1}$, $Z_{2}=\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}$ and $M=\prod_{\alpha=1}^{k}p_{\alpha}$. The value of $K$ depends on the choice of $\zeta=(\mathbf{s}_{1},\mathbf{s}_{2},\mathbf{t}_{1},\mathbf{t}_{2})\in\Gamma$ and can be calculated to be $K=2\prod_{\alpha=1}^{k}p_{\alpha}^{2}$. According to the newly proposed tighter upper bound, the proposed 2-D ZCACS becomes an optimal one when $p=2$.
VI Application in mMIMO
In this section, we discuss the applicability of the proposed 2-D array codes in mMIMO systems and their performances.
VI-A Technical Details of the Simulation
For the purpose of simulation, we have generated the steering matrix based on (19), where we have set $\lambda=3\times 10^{-1}$, $d_{x}=\lambda/2$, $d_{y}=\lambda/2$. The received signal is obtained by using (20). We have considered only full-rate STBCs in this simulation, which have been described in [29, Construction I]. To compare different precoding schemes, we have fixed the URA size of the particular simulation and varied the STBC configuration because it is impossible to fix both URA size and STBC size using 2-D array codes. We compared the performance of 2-D GCAS- and GCS-based precoders from Corollary 1 and Corollary 2, respectively, with that of the 2-D CCC-based precoding of [17], and 2-D GCAS-based precoding of [18, 19, 20]. The 2-D GCAPs from [21] have not been considered in this simulation, as $4\times 6$, $8\times 16$ and $64\times 54$- URA-based precoding matrices cannot be produced from the construction of [21]. To decode the signals in the simulation, we have used the maximum-likelihood (ML) technique. Finally, we have obtained the BER performance and plotted it against the signal-to-noise ratio (SNR) axis, which is computed in decibels (dB). We have used randomly generated binary phase shift keying (BPSK) data bits in the simulation, and in total, $4\times 10^{5}$ data bits are considered for each performance curve. The complex AWGN $w(t)$ used in the simulation is obtained using the formula:
$w(t)=\sigma\left(x+\iota y\right),$ (47)
where $\iota=\sqrt{-1}$; $x,y\in\mathcal{N}(0,1)$ are random variables, and $\mathcal{N}(0,1)$ denotes the standard normal distribution with mean 00, variance $1$, $\sigma^{2}=N_{0}/2$ with $N_{0}$ being the noise power, which is computed as $N_{0}=\left(\frac{\text{Signal Power}}{\text{SNR}}\right)$. Apart from the existing 2-D GCAS/CCC-based constructions, we have also compared random matrix (RM)-based and Zadoff-Chu (ZC) sequence-based precoders [37, 38], which are widely used in practice. For RM-based precoders, the matrix elements are drawn randomly from the set $\{+1,-1\}$. For the ZC-based precoder of URA size $P\times Q$, the $K$-precoding matrices are generated by using a length-$P$ ZC sequence $\mathbf{u}$ and a length-$Q$ ZC sequence $\mathbf{v}$, with the operation
$\mathbf{W}_{n}=\mathbf{u}_{n}\mathbf{v}_{n}^{H},\forall n=0,1,\dots,K-1;$ (48)
where $(\cdot)^{H}$ denotes the Hermitian operation, $\mathbf{u}_{i},\mathbf{v}_{j}$ denote the $i$-th and $j$-th cyclically shifted versions of $\mathbf{u}$ and $\mathbf{v}$, respectively [28]. In particular, in this simulation, for both RM-based and ZC-based precoders, we set $K=4$. To show the performance benefit of the precoders derived from our construction, we plot simulations for three different URA sizes, including both small and large URAs, i.e., $4\times 6$, $8\times 16$ and $64\times 54$-URA.
For the $4\times 6$ URA, the following precoders are used.
  1. (1)
    $4\times 4$-STBC is used for the precoder in [20]. The generating functions of $4$-ary $\mathbf{W}_{i}$’s used for [20] is given by $f=2(z_{1}z_{2}+z_{2}z_{3}+z_{3}z_{4})+1$.
  2. (2)
    $3\times 4$-STBC is used for the derived $6$-ary 2-D GCAS-based precoding matrices. The function for 2-D GCAS is given by (31), where $a_{\mathbf{s_{i}},\mathbf{t_{j}}}^{\gamma}=2v_{1}+2\gamma v_{1}+t_{j}v_{1}+2s_{i}v_{2}$, $a_{1}=3x_{1}+1$ and $b_{1},a_{2},b_{2}$ follows similarly.
For the $8\times 16$-URA, we use the following constructions.
  1. (1)
    $4\times 4$-STBC is used for precoders in [17, 18]. The generating functions of binary $\mathbf{W}_{i}$’s used for [18] is given by $f=(x_{1}x_{2}+x_{2}x_{3})+(y_{1}y_{2}+y_{2}y_{3}+y_{3}y_{4})+1$. For the case of [17] also, the same $\mathbf{W}_{i}$’s are used.
  2. (2)
    $8\times 8$-STBC is used for the case of [19] with the generating function of binary $\mathbf{W}_{i}$’s given by $f=(x_{1}x_{2}+x_{2}x_{3})+(y_{1}y_{2}+y_{3}y_{4})+1$.
  3. (3)
    $2\times 2$-Alamouti code [39] is used as STBC for the derived 2-D GCAS-based precoding matrices. The binary 2-D GCAS is constructed from Corollary 1, using the function given in (31), where $a_{\mathbf{s_{i}},\mathbf{t_{j}}}^{\gamma}=v_{1}v_{2}+v_{2}v_{3}+\gamma v_{1}+t_{j}v_{3}+s_{i}v_{4}$, $a_{1}=x_{1}x_{2}+1$, $a_{2},b_{1}$ and $b_{2}$ are derived using Lemma 6 and Lemma 7.
  4. (4)
    The derived binary GCS-based precoding matrix is generated from Corollary 2 with the function given by (31), where $a_{\mathbf{s_{i}},\mathbf{t_{j}}}^{\gamma}=v_{1}+\gamma v_{1}+t_{j}v_{1}+s_{i}v_{2}$, $a_{1}=x_{1}x_{2}+x_{2}x_{3}+1$. It uses the $8\times 8$-STBC model given in [29].
For the $64\times 54$-URA system, we compare the derived 2-D GCAS with the RM-based and ZC-based precoders only, because it is not feasible to construct such precoders and STBC combinations using [17, 18, 19, 20, 21]. The 2-D GCAS is derived as follows.
  1. (1)
    $3\times 4$-STBC is used for the derived 2-D GCAS-based precoder. The function for generating the $6$-ary 2-D GCAS is given by $a_{\mathbf{s_{i}},\mathbf{t_{j}}}^{\gamma}=2(v_{1}v_{2}+v_{2}v_{3})+2\gamma v_{1}+2t_{j}v_{3}+2s_{i}v_{4}$, $a_{1}=3(x_{1}x_{2}+x_{2}x_{3}+x_{3}x_{4}+x_{4}x_{5})+1$, $a_{1}=3(x_{1}x_{2}+x_{2}x_{3}+x_{3}x_{4}+x_{4}x_{5})+1$, and $b_{1},a_{2},b_{2}$ follows similarly.
VI-B Performance Analysis
In Fig. 3, the BER performances of the above-mentioned mMIMO-URA systems are shown with respect to varying SNR. Fig. 3a clearly shows the performance benefit of the derived 2-D arrays as precoding matrix in $4\times 6$-URA, over the existing constructions. The next figure shows the performance of precoders in larger URA size of $8\times 16$. Fig. 3b suggests that the derived 2-D GCAS and GCS perform as good as the 2-D CCC and 2-D GCAS-based precoding in [17, 18, 19], although the derived construction has benefits regarding the flexibility of the parameters. Besides that, the performances are better than widely used ZC sequence-based precoders. Note that, particularly for this example, from the point of view of complexity, the derived 2-D GCASs are more efficient than [19]. Because the precoder for [19] uses a large $8\times 8$-STBC with the same $8\times 16$-URA configuration increasing the transmission overhead, whereas the derived 2-D GCAS uses a small $2\times 2$-Alamouti scheme as STBC encoder without compromising the transmit diversity. Fig. 3c shows the performance benefit of the derived 2-D GCAS and GCS, when large-sized URAs are considered, $64\times 54$-URA for this example.
(a) BER comparison for $4\times 6$-URA
(b) BER comparison for $8\times 16$-URA
(c) BER comparison for $64\times 54$-URA
Figure 3: BER comparison for different URA sizes
VI-C Flexibility of System Parameters
The flock size of the 2-D array code determines the number of antennas to transmit STBC-encoded data. For the derived 2-D GCAS, flock size is of the form $p_{1}p_{2}\dots p_{k}$. Hence, it can support an arbitrary number of transmit antennas for STBC. Although increasing the transmit antennas for STBC encoders may result in higher complexity, here we only focus only on the flexibility of antenna size theoretically. The practical applicability of antenna size for STBC encoders would depend on the system in consideration. The same theory goes for URA sizes too. From the derived 2-D arrays, it appears that he URA sizes can also be made flexible. In Fig. 4, the supported URA sizes $P\times Q$ are plotted for different existing constructions [17, 18, 19, 21, 20] and we compare them with the derived 2-D GCAS and GCS. We have restricted the values of $P,Q$ in a finite range $[0,40]$ due to shortage of space. The figure shows that the proposed constructions support many such URA sizes that cannot be supported by the existing constructions, as also shown in the next example.
Example 5
Consider the transmission through $18$-antenna STBC encoder in the mMIMO with URA size $4\times 18$. This can be achieved only by the GCS derived from Corollary 2, but not by [17, 18]. Similarly, the transmission through $3$-antenna STBC encoder with URA size $4\times 6$ can be achieved only by the 2-D GCAS derived in Corollary 1, but not by [17, 18].
Figure 4: Plot of URA sizes for different constructions
In the next section, we compare the proposed construction with the existing ones regarding their parameter flexibility.
VII Comparison with Existing Works
In TABLE II, we compare the direct constructions of 2-D array codes used for the mMIMO system in [17, 18, 19, 20, 21] with our derived 2-D GCAS and GCS, based on their parameters. Also, in TABLE III, we compare 2-D ZCACSs in [23, 40, 41, 42] with our proposed construction of 2-D ZCACS. It can be seen from TABLE II and III that our proposed constructions provide a flexible set of parameters that are unreported and also can be formulated in explicit mathematical forms.
TABLE II: Comparison table URA-STBC configuration using different 2-D arrays
Construction Array/URA size Flock size STBC required (full-rate) Constraints
[17] $2^{m}\times 2^{n}$ $2^{k}$ $2^{k}\times 2^{k}$ $m,n,k\geq 1,m,n\geq k$
[18] $p^{m_{1}}\times p^{m_{2}}$ $p^{k}$ $p^{k}\times N$ $p$ is a prime, $m_{1},m_{2},k\geq 1$, $m_{1}+m_{2}\geq k$, $p^{k}\leq\mu(N)$
[19] $b_{1}^{m_{1}}\times b_{2}^{m_{2}}$ $N_{1}^{k_{1}}N_{2}^{k_{2}}$ $N_{1}^{k_{1}}N_{2}^{k_{2}}\times N$ $N_{1}\geq b_{1}$, $N_{1}\geq b_{1}$, $m_{1},m_{2}\geq 1$, $b_{1},b_{2}\geq 2$, $k_{1}\leq m_{1}$, $k_{2}\leq m_{2}$, $N_{1}^{k_{1}}N_{2}^{k_{2}}\leq\mu(N)$
[20] $2^{n}\times N$ $2^{k+1}$ $2^{k+1}\times 2^{k+1}$ $N\in\mathbb{Z}^{+},n\geq 2,k\in\mathbb{Z}^{+}\cup\{0\}$
[21] $2^{m_{1}}\times 10\cdot 2^{m_{2}-4}$, $2^{m_{3}}\times 26\cdot 2^{m_{4}-5}$ $2$ $2\times 2$ $m_{1},m_{3}\geq 1,m_{2}\geq 5,m_{4}\geq 6$
Proposed, Corollary 1 $2^{m+1}\times 2p_{1}^{m_{1}-1}\cdots p_{k}^{m_{k}-1}$ $p_{1}p_{2}\cdots p_{k}$ $p_{1}p_{2}\cdots p_{k}\times N$ $m\geq 1,m_{i}\geq 2,\forall i,k\in\mathbb{N},p_{1}p_{2}\cdots p_{k}\leq\mu(N)$
Proposed, Corollary 2 $2^{m+1}\times 2p_{1}^{m_{1}}\cdots p_{k}^{m_{k}}$ $2p_{1}^{m_{1}}\cdots p_{k}^{m_{k}}$ $2p_{1}^{m_{1}}\cdots p_{k}^{m_{k}}\times N$ $m\geq 1,m_{i}\geq 2,\forall i,k\in\mathbb{N},2p_{1}^{m_{1}}\cdots p_{k}^{m_{k}}\leq\mu\left(N\right)$
[…]
TABLE III: Comparison table for 2-D ZCACS
Construction Phase Array size Rectangular ZCZ size Set size Flock size Constraints Based on
[23] $\{1,-1,0\}$ $L_{1}\times(L_{2}+r+1)$ $Z_{1}\times Z_{2}$ $K=K^{\prime}r$ 2 $L_{1},L_{2}\geq 1$; $Z_{1},Z_{2}\geq 1$; $rZ_{2}\leq Z^{\prime}_{2}$; $K\geq 1$; $L_{1}\times L_{2}$ and $Z^{\prime}_{1}\times Z^{\prime}_{2}$ are the size and ZCZ size of seed 2-D ZCACS with set size $K^{\prime}$ Matrix operation
[40] $q$ $L_{1}\times L_{2}$ $Z_{1}\times Z_{2}$ $K$ $M$ $q=lcm(q_{\mathbf{V}_{K}},q_{\mathbf{U}_{K}},q_{\mathbf{U}_{M}})$; $q\geq 2$, $L_{1}=L_{2}=K$; $Z_{1}=M,Z_{2}=K$, $K=MP$ Butson-type Hadamard (BH) matrices
[41] $q$ $R_{1}N_{1}\times R_{2}N_{2}$ $N_{1}\times N_{2}$ $R_{1}R_{2}N_{1}N_{2}$ $N_{1}N_{2}$ $R_{1}\geq 1,R_{2}\geq 2$, $N_{1}=p_{1}^{m_{1}}\cdots p_{k_{1}}^{m_{k_{1}}}$, $N_{2}=q_{1}^{n_{1}}\cdots q_{k_{2}}^{n_{k_{2}}}$, $p_{i},q_{i}$ are prime, $m_{i},n_{i}\geq 1$, $p_{i}\mid q$, $q_{j}\mid q,\forall i,j$ EBF
[42] $q$ $p^{m_{1}}\times p^{m_{2}}$ $p^{m_{1}-s_{1}}\times p^{m_{2}-s_{2}}$ $p^{k+s+1}$ $p^{k+1}$ $p$ is prime, $p\mid q$, $k+s<m-1$, $k_{1}+s_{1}\leq m_{1}$, $k_{2}+s_{2}\leq m_{2}$ EBF
Proposed, Theorem 1 $q$ $L_{1}\times L_{2}$ $Z_{1}\times Z_{2}$ $2p_{1}^{2}p_{2}^{2}\dots p_{k}^{2}$ $p_{1}p_{2}\dots p_{k}$ $L_{1}=2^{m}p$, $L_{2}=2\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}$; $Z_{1}=2^{m+1}$, $Z_{2}=\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}$; $p_{\alpha}$ is prime, $\forall\alpha$; $p$ is prime; $m\geq 1$, $m_{\alpha}\geq 2$, $\forall\alpha$; $2\mid q,p\mid q,p_{\alpha}\mid q$, $\forall\alpha$ EBF
VIII Conclusion
In this paper, we first constructed IGC code set as a foundation, and using it 2-D ZCAC and 2-D ZCACS with flexible array sizes have been proposed. The bounds of row and column sequence PMEPRs have been investigated and shown to have advantages over existing constructions. We have also proposed a new upper bound for the set size of 2-D ZCACS. The proposed construction is optimal with respect to the theoretical upper bound for special cases. In addition to that, 2-D GCAS and GCS are derived from the proposed 2-D ZCAC and applied as precoding matrices in the mMIMO systems with URA-STBC for omnidirectional transmission. Simulation results show that the derived 2-D arrays have better BER performance than the ones in practice. Also, they offer a wide range of flexibility in terms of URA size and the number of antennas for STBC encoding. For future research, an interesting work could be to construct 2-D ZCACS with more flexibility.
Appendix A Proof of Lemma 5
We state two lemmas before the proof of Lemma 5.
Lemma 9 ([43])
For $\mathbf{t}_{1},\mathbf{t}_{2}\in\mathbb{Z}_{p_{1}}\times\mathbb{Z}_{p_{2}}\times\dots\times\mathbb{Z}_{p_{k}}$, we have
$\begin{split}&\sum_{\bm{\gamma\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}}C\left(\psi(R_{\mathbf{t}_{1}}^{\bm{\gamma}}),\psi(R_{\mathbf{t}_{2}}^{\bm{\gamma}})\right)(\tau)\\ &=\begin{cases}\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}},&\mathbf{t}_{1}=\mathbf{t}_{2},\tau=0;\\ 0,&\mathbf{t}_{1}=\mathbf{t}_{2},0<\mid\tau\mid<\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1};\\ 0,&\mathbf{t}_{1}\neq\mathbf{t}_{2},\mid\tau\mid<\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}.\end{cases}\end{split}$ (49)
This result follows from the construction of MOGCS in [43].
Lemma 10
For $\mathbf{s}_{1},\mathbf{s}_{2}\in\mathbb{Z}_{p_{1}}\times\mathbb{Z}_{p_{2}}\times\dots\times\mathbb{Z}_{p_{k}}$, we have
$C\left(\psi(T_{\mathbf{s}_{1}}),\psi(T_{\mathbf{s}_{2}})\right)(0)=\begin{cases}\prod_{\alpha=1}^{k}p_{\alpha},&\mathbf{s}_{1}=\mathbf{s}_{2};\\ 0,&\mathbf{s}_{1}\neq\mathbf{s}_{2}.\end{cases}$ (50)
Proof:
We let $\mathbf{s}_{i}=(s^{i}_{1},s^{i}_{2},\dots,s^{i}_{k})$ for $i=1,2$. From the definition of $T_{\mathbf{s}}$, we have
$\begin{split}&C\left(\psi(T_{\mathbf{s}_{1}}),\psi(T_{\mathbf{s}_{2}})\right)(0)\\ &=\sum_{v^{\prime}_{p_{1}}=0}^{p_{1}-1}\sum_{v^{\prime}_{p_{2}}=0}^{p_{2}-1}\dots\sum_{v^{\prime}_{p_{k}}=0}^{p_{k}-1}\omega_{q}^{\sum_{\alpha=1}^{k}\frac{q}{p_{\alpha}}v^{\prime}_{p_{\alpha}}(s^{1}_{\alpha}-s^{2}_{\alpha})}\\ &=\sum_{v^{\prime}_{p_{1}}=0}^{p_{1}-1}\omega_{p_{1}}^{v^{\prime}_{p_{1}}(s^{1}_{1}-s^{2}_{1})}\times\dots\times\sum_{v^{\prime}_{p_{k}}=0}^{p_{k}-1}\omega_{p_{k}}^{v^{\prime}_{p_{k}}(s^{1}_{k}-s^{2}_{k})}.\end{split}$ (51)
When $\mathbf{s}_{1}=\mathbf{s}_{2}$, then $s^{1}_{\alpha}=s^{2}_{\alpha},\forall\alpha$. So, we have
$\sum_{v^{\prime}_{p_{\alpha}}=0}^{p_{\alpha}-1}\omega_{p_{\alpha}}^{v^{\prime}_{p_{\alpha}}(s^{1}_{\alpha}-s^{2}_{\alpha})}=p_{\alpha},\forall\alpha.$ (52)
From (51), (52) for $\mathbf{s}_{1}=\mathbf{s}_{2}$ we have
$C\left(\psi(T_{\mathbf{s}_{1}}),\psi(T_{\mathbf{s}_{2}})\right)(0)=\prod_{\alpha=1}^{k}p_{\alpha}.$ (53)
On the other hand, when $\mathbf{s}_{1}\neq\mathbf{s}_{2}$, then $s^{1}_{\delta}\neq s^{2}_{\delta}$ for some $\delta\in\{1,2,\dots,k\}$. So, we have
$\sum_{v^{\prime}_{p_{\delta}}=0}^{p_{\delta}-1}\omega_{p_{\delta}}^{v^{\prime}_{p_{\delta}}(s^{1}_{\delta}-s^{2}_{\delta})}=0,$ (54)
because $\{\omega_{p_{\delta}}^{v^{\prime}_{p_{\delta}}(s^{1}_{\delta}-s^{2}_{\delta})}:v^{\prime}_{p_{\delta}}=0,1,\dots,p_{\delta}-1\}$ are all the roots of the polynomial $z^{p_{\delta}}-1$. Hence, from (51) and (54), for $\mathbf{s}_{1}\neq\mathbf{s}_{2}$ we have the result. ∎
Now we come to the main proof of Lemma 5. Note that
$\psi(a_{\mathbf{s},\mathbf{t}}^{\bm{\gamma}})=\psi(T_{\mathbf{s}})\otimes\psi(R_{\mathbf{t}}^{\bm{\gamma}}),$ (55)
where $\otimes$ denotes the Kronecker product. Then we have
$\begin{split}&\sum_{\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}C\left(\psi(a_{\mathbf{s}_{1},\mathbf{t}_{1}}^{\bm{\gamma}}),\psi(a_{\mathbf{s}_{2},\mathbf{t}_{2}}^{\bm{\gamma}})\right)(\tau)\\ &=\bigg{[}C\left(\psi(T_{\mathbf{s}_{1}}),\psi(T_{\mathbf{s}_{2}})\right)\left(\left\lfloor\frac{\tau}{N}\right\rfloor\right)\\ &\quad\times\sum_{\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}C\left(\psi(R_{\mathbf{t}_{1}}^{\bm{\gamma}}),\psi(R_{\mathbf{t}_{2}}^{\bm{\gamma}})\right)\left(\tau\mod N\right)\bigg{]}\\ &\quad+\bigg{[}\Delta_{N}C\left(\psi(T_{\mathbf{s}_{1}}),\psi(T_{\mathbf{s}_{2}})\right)\left(\left\lfloor\frac{\tau}{N}\right\rfloor+1\right)\\ &\quad\times\sum_{\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}C\left(\psi(R_{\mathbf{t}_{1}}^{\bm{\gamma}}),\psi(R_{\mathbf{t}_{2}}^{\bm{\gamma}})\right)\left(\tau\mod N-N\right)\bigg{]},\end{split}$ (56)
with $N=\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}$, because for two pairs of sequences $(\mathbf{a}_{1},\mathbf{a}_{2})$, $(\mathbf{b}_{1},\mathbf{b}_{2})$ and $\tau\geq 0$, we have
$\begin{split}&C(\mathbf{a}_{1}\otimes\mathbf{b}_{1},\mathbf{a}_{2}\otimes\mathbf{b}_{2})(\tau)\\ &=C(\mathbf{a}_{1},\mathbf{a}_{2})\left(\left\lfloor\frac{\tau}{N}\right\rfloor\right)C(\mathbf{b}_{1},\mathbf{b}_{2})(\tau\!\!\!\!\mod N)\\ &\quad+\!\Delta_{N}C(\mathbf{a}_{1},\mathbf{a}_{2})\!\left(\left\lfloor\frac{\tau}{N}\right\rfloor\!\!+\!\!1\right)\!C(\mathbf{b}_{1},\mathbf{b}_{2})(\tau\!\!\!\!\!\mod N\!\!-\!\!N),\end{split}$ (57)
where each $\mathbf{a}_{i}$ has length $M$ and each $\mathbf{b}_{i}$ has length $N$, and
$\Delta_{N}=\begin{cases}0,&\tau\mod N=0;\\ 1,&\text{otherwise}.\end{cases}$ (58)
We need to show that
$\begin{split}&\sum_{\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}C\left(\psi(a_{\mathbf{s}_{1},\mathbf{t}_{1}}^{\bm{\gamma}}),\psi(a_{\mathbf{s}_{2},\mathbf{t}_{2}}^{\bm{\gamma}})\right)(\tau)\\ &=\begin{cases}\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}+1},&\left(\mathbf{s}_{1},\mathbf{t}_{1}\right)=\left(\mathbf{s}_{2},\mathbf{t}_{2}\right),\tau=0;\\ 0,&\left(\mathbf{s}_{1},\mathbf{t}_{1}\right)=\left(\mathbf{s}_{2},\mathbf{t}_{2}\right),\\ &0<\tau<\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1};\\ 0,&\mathbf{s}_{1}\neq\mathbf{s}_{2},\mathbf{t}_{1}=\mathbf{t}_{2},\tau<\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}.\\ \end{cases}\end{split}$ (59)
Here we considered only $\tau\geq 0$, because $C(\mathbf{a},\mathbf{b})(-\tau)=C^{*}(\mathbf{b},\mathbf{a})(\tau)$. The rest of the proof is split into three cases.
  1. Case I:
    Let $\tau=0$. Then $\left\lfloor\frac{\tau}{N}\right\rfloor=0$, $\tau\mod N=0$ and $\Delta_{N}=0$. Using Lemma 9, Lemma 10, and (56), we have
    $\begin{split}&\sum_{\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}C\left(\psi(a_{\mathbf{s}_{1},\mathbf{t}_{1}}^{\bm{\gamma}}),\psi(a_{\mathbf{s}_{2},\mathbf{t}_{2}}^{\bm{\gamma}})\right)(\tau)\\ &=\begin{cases}\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}+1},&\mathbf{s}_{1}=\mathbf{s}_{2},\mathbf{t}_{1}=\mathbf{t}_{2};\\ 0,&\mathbf{s}_{1}\neq\mathbf{s}_{2},\mathbf{t}_{1}=\mathbf{t}_{2};\\ 0,&\mathbf{t}_{1}\neq\mathbf{t}_{2}.\end{cases}\end{split}$ (60)
  2. Case II:
    Let $0<\tau<\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}$. In this case $\left\lfloor\frac{\tau}{N}\right\rfloor=0$, $0<\tau\mod N<\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}$, $-\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}<\tau\mod N-N<0$ and $\Delta_{N}=1$. Using Lemma 9, we have
    $\begin{split}\sum_{\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}C\left(\psi(R_{\mathbf{t}_{1}}^{\bm{\gamma}}),\psi(R_{\mathbf{t}_{2}}^{\bm{\gamma}})\right)\left(\tau\mod N\right)=&0,\\ \sum_{\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}\!\!\!\!\!C\left(\psi(R_{\mathbf{t}_{1}}^{\bm{\gamma}}),\psi(R_{\mathbf{t}_{2}}^{\bm{\gamma}})\right)\left(\tau\mod N-N\right)=&0,\end{split}$ (61)
    for all values of $\mathbf{t}_{1}$ and $\mathbf{t}_{2}$. Then, from (56) and (61), for all choices of $\left(\mathbf{s}_{1},\mathbf{t}_{1}\right)$ and $\left(\mathbf{s}_{2},\mathbf{t}_{2}\right)$, we have
    $\begin{split}\sum_{\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}\!\!\!\!\!C\left(\psi(a_{\mathbf{s}_{1},\mathbf{t}_{1}}^{\bm{\gamma}}),\psi(a_{\mathbf{s}_{2},\mathbf{t}_{2}}^{\bm{\gamma}})\right)(\tau)=0.\end{split}$ (62)
  3. Case III:
    Let $\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}\leq\tau<\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}$ and $\mathbf{t}_{1}\neq\mathbf{t}_{2}$. In this case, we have two sub-cases.
    1. Sub-case 1:
      Let $\tau=\beta\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}$ for some $\beta\in\{1,2,\dots,\prod_{\alpha=1}^{k}p_{\alpha}-1\}$. In that case $\left\lfloor\frac{\tau}{N}\right\rfloor=\beta$, $\tau\mod N=0$ and $\Delta_{N}=0$. Then, form Lemma 9, we have
      $\sum_{\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}C\left(\psi(R_{\mathbf{t}_{1}}^{\bm{\gamma}}),\psi(R_{\mathbf{t}_{2}}^{\bm{\gamma}})\right)\left(\tau\!\!\!\!\mod N\right)=0.$ (63)
    2. Sub-case 2:
      Let $\beta\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}<\tau<(\beta+1)\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}$ for some $\beta\in\{1,2,\dots,\prod_{\alpha=1}^{k}p_{\alpha}-1\}$. Then $\left\lfloor\frac{\tau}{N}\right\rfloor=\beta$, $0<\tau\mod N<\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}$, and $-\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}<\tau\mod N-N<0$ with $\Delta_{N}=1$. Then using Lemma 9, we have
      $\begin{split}\sum_{\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}\!\!\!\!\!C\left(\psi(R_{\mathbf{t}_{1}}^{\bm{\gamma}}),\psi(R_{\mathbf{t}_{2}}^{\bm{\gamma}})\right)\left(\tau\!\!\!\!\mod N\right)=&0,\\ \sum_{\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}\!\!\!\!\!\!\!\!\!\!C\left(\psi(R_{\mathbf{t}_{1}}^{\bm{\gamma}}),\psi(R_{\mathbf{t}_{2}}^{\bm{\gamma}})\right)\left(\tau\!\!\!\!\mod N-N\right)=&0.\end{split}$ (64)
Hence, the lemma is proved.
Appendix B Proof of Theorem 1
To prove the theorem, we need to show that
$\begin{split}&\sum_{\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}C\left(\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{1}},\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{2}}\right)(\tau_{1},\tau_{2})\\ &=\begin{cases}L_{1}L_{2}\prod_{\alpha=1}^{k}p_{\alpha},&\bm{\zeta}_{1}=\bm{\zeta}_{2},(\tau_{1},\tau_{2})=(0,0);\\ 0,&\bm{\zeta}_{1}=\bm{\zeta}_{2},(\tau_{1},\tau_{2})\neq(0,0);\\ &\left|\tau_{1}\right|<2^{m+1},\left|\tau_{2}\right|<\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1};\\ 0,&\bm{\zeta}_{1}\neq\bm{\zeta}_{2},\left|\tau_{1}\right|<2^{m+1},\\ &\left|\tau_{2}\right|<\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1},\\ \end{cases}\end{split}$ (65)
where $L_{1}=2^{m}p$ $L_{2}=2\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}$. It is straightforward to verify that the number of $\bm{\zeta}\in\Gamma$ is $K=2p_{1}^{2}p_{2}^{2}\dots p_{k}^{2}$. Also, for the trivial case $\bm{\zeta}_{1}=\bm{\zeta}_{2},(\tau_{1},\tau_{2})=(0,0)$, we have
$\begin{split}\sum_{\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}C\left(\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{1}},\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{2}}\right)(\tau_{1},\tau_{2})=2^{m+1}p\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}+1},\end{split}$ (66)
To prove the rest of the theorem, we shall consider only $\tau_{1}\geq 0$, and $0\leq\left|\tau_{2}\right|<\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}$, because for any two 2-D arrays $\mathbf{X}_{1}$, $\mathbf{X}_{2}$, we have $C(\mathbf{X}_{1},\mathbf{X}_{2})(-\tau_{1},\tau_{2})=C(\mathbf{X}_{2},\mathbf{X}_{1})^{*}(\tau_{1},-\tau_{2})$. For a fixed $\bm{\gamma}$, we have
$\begin{split}&C\left(\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{1}},\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{2}}\right)(\tau_{1},\tau_{2})\\ &=\sum_{i=0}^{L_{1}-\tau_{1}-1}\sum_{j=0}^{L_{2}-\tau_{2}-1}\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{1}}_{i,j}{\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{2}}_{i^{\prime},j^{\prime}}}^{*}\\ &=\sum_{i=0}^{L_{1}-\tau_{1}-1}\sum_{j=0}^{L_{2}-\tau_{2}-1}\omega_{q}^{F_{\vec{i}}^{\bm{\gamma},\bm{\zeta}_{1}}(\vec{j})}\omega_{q}^{-F_{\vec{i^{\prime}}}^{\bm{\gamma},\bm{\zeta}_{2}}(\vec{j^{\prime}})},\end{split}$ (67)
where $i^{\prime}=i+\tau_{1}$, $j^{\prime}=j+\tau_{2}$. Let $1+\sum_{\alpha=1}^{k}m_{\alpha}=\epsilon$ and $m+1=\vartheta$. We use the following notations.
  1. (1)
    We let $\vec{i}=(\vec{i}_{\delta},i_{\vartheta})$ and $\vec{i^{\prime}}=(\vec{i}^{\prime}_{\delta},i^{\prime}_{\vartheta})$, where $\vec{i}_{\delta},\vec{i}^{\prime}_{\delta}$ are the vectors with values in $\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}^{m_{\alpha}-1}\times\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}$ which are truncated from $\vec{i}$ and $\vec{i^{\prime}}$, respectively. Whereas, $i_{\vartheta},i^{\prime}_{\vartheta}$ are the last elements of $\vec{i},\vec{i}^{\prime}$ corresponding to $\mathbb{Z}_{2}$.
  2. (2)
    Similarly, we define $\vec{j}=(\vec{j}_{\xi},j_{\epsilon})$ and $\vec{j}^{\prime}=(\vec{j}^{\prime}_{\xi},j^{\prime}_{\epsilon})$, where $\vec{j}_{\xi},\vec{j}^{\prime}_{\xi}\in\mathbb{Z}_{2}^{m}$ and $\vec{j}_{\epsilon},\vec{j}^{\prime}_{\epsilon}\in\mathbb{Z}_{p}$.
  3. (3)
    We define $i_{\delta},i^{\prime}_{\delta}$ to be the decimal numbers derived from the radix-representations $\vec{i}_{\delta},\vec{i}^{\prime}_{\delta}$, respectively. Similarly, statement holds for $j_{\xi},j^{\prime}_{\xi}$.
Then, expanding the function $F_{\vec{d}^{\prime}}^{\bm{\gamma},\bm{\zeta}}(\vec{x})$, we have
$\begin{split}&\omega_{q}^{F_{\vec{i}}^{\bm{\gamma},\bm{\zeta}_{1}}(\vec{j})}\omega_{q}^{-F_{\vec{i}^{\prime}}^{\bm{\gamma},\bm{\zeta}_{2}}(\vec{j}^{\prime})}\\ &=\omega_{q}^{(1-j_{\epsilon})\left\{a_{\mathbf{s}_{1}^{1},\mathbf{t}_{1}^{1}}^{\bm{\gamma}}(\vec{j}_{\xi})+a_{1}(\vec{i}_{\delta})W_{1}(i_{\vartheta})+a_{2}(\vec{i}_{\delta})W_{2}(i_{\vartheta})\right\}}\\ &\quad\times\omega_{q}^{j_{\epsilon}\left\{a_{\mathbf{s}_{2}^{1},\mathbf{t}_{2}^{1}}^{\bm{\gamma}}(\vec{j}_{\xi})+b_{1}(\vec{i}_{\delta})W_{1}(i_{\vartheta})+b_{2}(\vec{i}_{\delta})W_{2}(i_{\vartheta})\right\}}\\ &\quad\times\omega_{q}^{-(1-j^{\prime}_{\epsilon})\left\{a_{\mathbf{s}_{1}^{2},\mathbf{t}_{1}^{2}}^{\bm{\gamma}}(\vec{j}^{\prime}_{\xi})+a_{1}(\vec{i}^{\prime}_{\delta})W_{1}(i^{\prime}_{\vartheta})+a_{2}(\vec{i}^{\prime}_{\delta})W_{2}(i^{\prime}_{\vartheta})\right\}}\\ &\quad\times\omega_{q}^{-j^{\prime}_{\epsilon}\left\{a_{\mathbf{s}_{2}^{2},\mathbf{t}_{2}^{2}}^{\bm{\gamma}}(\vec{j}^{\prime}_{\xi})+b_{1}(\vec{i}^{\prime}_{\delta})W_{1}(i^{\prime}_{\vartheta})+b_{2}(\vec{i}^{\prime}_{\delta})W_{2}(i^{\prime}_{\vartheta})\right\}},\end{split}$ (68)
Now, we split the rest of the proof into three cases.
  1. Case I:
    Let $0\leq\tau_{1}<2^{m+1}$ and $0<\tau_{2}<\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}$. Then define
    $\begin{split}S_{1}=&\sum_{j=0}^{\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}-\tau_{2}-1}\omega_{q}^{F_{\vec{i}}^{\bm{\gamma},\bm{\zeta}_{1}}(\vec{j})}\omega_{q}^{-F_{\vec{i}^{\prime}}^{\bm{\gamma},\bm{\zeta}_{2}}(\vec{j}^{\prime})},\\ S_{2}=&\sum_{j=\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}-\tau_{2}}^{\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}-1}\omega_{q}^{F_{\vec{i}}^{\bm{\gamma},\bm{\zeta}_{1}}(\vec{j})}\omega_{q}^{-F_{\vec{i}^{\prime}}^{\bm{\gamma},\bm{\zeta}_{2}}(\vec{j}^{\prime})},\\ S_{3}=&\sum_{j=\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}}^{2\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}-\tau_{2}-1}\omega_{q}^{F_{\vec{i}}^{\bm{\gamma},\bm{\zeta}_{1}}(\vec{j})}\omega_{q}^{-F_{\vec{i}^{\prime}}^{\bm{\gamma},\bm{\zeta}_{2}}(\vec{j}^{\prime})}.\\ \end{split}$ (69)
    Then, we can write the term $C\left(\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{1}},\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{2}}\right)(\tau_{1},\tau_{2})$ in the following form:
    $C\left(\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{1}},\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{2}}\right)(\tau_{1},\tau_{2})=\sum_{i=0}^{L_{1}-\tau_{1}-1}\bigg{[}S_{1}+S_{2}+S_{3}\bigg{]}.$ (70)
    1. (1)
      For $S_{1}$, we have $10\leq j\leq\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}-\tau_{2}-$ 1. This implies that $j_{\epsilon}=j^{\prime}_{\epsilon}=0$ as $\tau_{2}<\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}$. Similarly, we can find that $10\leq j_{\xi}\leq\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}-\tau_{2}-$ 1, $\tau_{2}\leq j^{\prime}_{\xi}\leq\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}-1$. So, using (68), we get
      $\begin{split}S_{1}=&\omega_{q}^{a_{1}(\vec{i}_{\delta})W_{1}(i_{\vartheta})+a_{2}(\vec{i}_{\delta})W_{2}(i_{\vartheta})}\omega_{q}^{-\left(a_{1}(\vec{i}^{\prime}_{\delta})W_{1}(i^{\prime}_{\vartheta})+a_{2}(\vec{i}^{\prime}_{\delta})W_{2}(i^{\prime}_{\vartheta})\right)}\\ &\times\left(\sum_{j_{\xi}=0}^{\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}-\tau_{2}-1}\omega_{q}^{a_{\mathbf{s}_{1}^{1},\mathbf{t}_{1}^{1}}^{\bm{\gamma}}(\vec{j}_{\xi})-a_{\mathbf{s}_{1}^{2},\mathbf{t}_{1}^{2}}^{\bm{\gamma}}(\vec{j}^{\prime}_{\xi})}\right)\\ \end{split}$ (71)
      Then, whether or not $\mathbf{s}_{1}^{1}=\mathbf{s}_{1}^{2}$ or $\mathbf{t}_{1}^{1}=\mathbf{t}_{1}^{2}$, i.e., irrespective of $\bm{\zeta}_{1}=\bm{\zeta}_{2}$ or $\bm{\zeta}_{1}\neq\bm{\zeta}_{2}$, from Lemma 5, we can conclude
      $\sum_{\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}\sum_{j_{\xi}=0}^{\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}-\tau_{2}-1}\omega_{q}^{a_{\mathbf{s}_{1}^{1},\mathbf{t}_{1}^{1}}^{\bm{\gamma}}(\vec{j}_{\xi})-a_{\mathbf{s}_{1}^{2},\mathbf{t}_{1}^{2}}^{\bm{\gamma}}(\vec{j}^{\prime}_{\xi})}=0,$ (72)
      because $0\lneqq\tau_{2}<\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}$.
    2. (2)
      For $S_{2}$, using a similar argument, we get
      $\begin{split}\sum_{\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}\sum_{j_{\xi}=\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}-\tau_{2}}^{\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}-1}\omega_{q}^{a_{\mathbf{s}_{1}^{1},\mathbf{t}_{1}^{1}}^{\bm{\gamma}}(\vec{j}_{\xi})-a_{\mathbf{s}_{2}^{2},\mathbf{t}_{2}^{2}}^{\bm{\gamma}}(\vec{j}^{\prime}_{\xi})}=0.\end{split}$ (73)
    3. (3)
      For $S_{3}$, similarly we have
      $\begin{split}\sum_{\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}\sum_{j_{\xi}=\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}}^{2\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}-\tau_{2}-1}\omega_{q}^{F_{\vec{i}}^{\bm{\gamma},\bm{\zeta}_{1}}(\vec{j}_{\xi})}\omega_{q}^{-F_{\vec{i}^{\prime}}^{\bm{\gamma},\bm{\zeta}_{2}}(\vec{j}^{\prime}_{\xi})}=0.\end{split}$ (74)
    Now, combining (72), (73), (74), and (70), we have
    $\begin{split}\sum_{\bm{\gamma}\in\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}C\left(\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{1}},\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{2}}\right)(\tau_{1},\tau_{2})=0.\end{split}$ (75)
  2. Case II:
    Let $\tau_{2}=0$, and $0<\tau_{1}<2^{m+1}$. Then, we have
    $\begin{split}C\left(\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{1}},\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{2}}\right)(\tau_{1},\tau_{2})=\sum_{i=0}^{L_{1}-\tau_{1}-1}\bigg{[}T_{1}+T_{2}\bigg{]},\\ \end{split}$ (76)
    where
    $\begin{split}T_{1}=&\sum_{j=0}^{\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}-1}\omega_{q}^{F_{\vec{i}}^{\bm{\gamma},\bm{\zeta}_{1}}(\vec{j})}\omega_{q}^{-F_{\vec{i}^{\prime}}^{\bm{\gamma},\bm{\zeta}_{2}}(\vec{j})},\\ T_{2}=&\sum_{j=\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}}^{2\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}-1}\omega_{q}^{F_{\vec{i}}^{\bm{\gamma},\bm{\zeta}_{1}}(\vec{j})}\omega_{q}^{-F_{\vec{i}^{\prime}}^{\bm{\gamma},\bm{\zeta}_{2}}(\vec{j})}.\end{split}$ (77)
    As $\tau_{2}=0$, this implies that for $T_{1}$ we have $j_{\epsilon}=0$, and for $T_{2}$ we have $j_{\epsilon}=1$. Hence, from (68), we have
    $\begin{split}T_{1}=&\omega_{q}^{a_{1}(\vec{i}_{\delta})W_{1}(i_{\vartheta})+a_{2}(\vec{i}_{\delta})W_{2}(i_{\vartheta})}\omega_{q}^{-\left(a_{1}(\vec{i}^{\prime}_{\delta})W_{1}(i^{\prime}_{\vartheta})+a_{2}(\vec{i}^{\prime}_{\delta})W_{2}(i^{\prime}_{\vartheta})\right)}\\ &\times\left(\sum_{j_{\xi}=0}^{\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}-1}\omega_{q}^{a_{\mathbf{s}_{1}^{1},\mathbf{t}_{1}^{1}}^{\bm{\gamma}}(\vec{j}_{\xi})-a_{\mathbf{s}_{1}^{2},\mathbf{t}_{1}^{2}}^{\bm{\gamma}}(\vec{j}_{\xi})}\right),\\ \end{split}$ (78)
    and
    $\begin{split}T_{2}=&\omega_{q}^{b_{1}(\vec{i}_{\delta})W_{1}(i_{\vartheta})+b_{2}(\vec{i}_{\delta})W_{2}(i_{\vartheta})}\omega_{q}^{-\left(b_{1}(\vec{i}^{\prime}_{\delta})W_{1}(i^{\prime}_{\vartheta})+b_{2}(\vec{i}^{\prime}_{\delta})W_{2}(i^{\prime}_{\vartheta})\right)}\\ &\times\left(\sum_{j_{\xi}=\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}}^{2\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}-1}\omega_{q}^{a_{\mathbf{s}_{2}^{1},\mathbf{t}_{2}^{1}}^{\bm{\gamma}}(\vec{j}_{\xi})-a_{\mathbf{s}_{2}^{2},\mathbf{t}_{2}^{2}}^{\bm{\gamma}}(\vec{j}_{\xi})}\right).\\ \end{split}$ (79)
    Now, we have two sub-cases.
    1. Sub-case 1:
      Let $\bm{\zeta}_{1}=\bm{\zeta}_{2}$. In that case, we have
      $\begin{split}\sum_{j_{\xi}=0}^{\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}-1}\omega_{q}^{a_{\mathbf{s}_{1}^{1},\mathbf{t}_{1}^{1}}^{\bm{\gamma}}(\vec{j}_{\xi})-a_{\mathbf{s}_{1}^{2},\mathbf{t}_{1}^{2}}^{\bm{\gamma}}(\vec{j}_{\xi})}=&\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}},\\ \sum_{j_{\xi}=\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}}^{2\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}-1}\omega_{q}^{a_{\mathbf{s}_{2}^{1},\mathbf{t}_{2}^{1}}^{\bm{\gamma}}(\vec{j}_{\xi})-a_{\mathbf{s}_{2}^{2},\mathbf{t}_{2}^{2}}^{\bm{\gamma}}(\vec{j}_{\xi})}=&\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}.\\ \end{split}$ (80)
      Also, using the property of $W_{1}$ and $W_{2}$, we can find that
      $\begin{split}W_{1}(i_{\vartheta})=0\Leftrightarrow&~{}W_{2}(i_{\vartheta})=1,\\ W_{1}(i_{\vartheta})=1\Leftrightarrow&~{}W_{2}(i_{\vartheta})=0.\\ \end{split}$ (81)
      Hence, without loss of generality, from the fact that $0<\tau_{1}\lneqq 2^{m+1}$ and from (78) and (79), we get
      $\begin{split}&\sum_{i=0}^{L_{1}-\tau_{1}-1}\bigg{[}T_{1}+T_{2}\bigg{]}\\ &=\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}}\sum_{i_{\delta}=0}^{L_{1}-\tau_{1}-1}\bigg{[}\left(\omega_{q}^{a_{1}(\vec{i}_{\delta})-a_{1}(\vec{i}^{\prime}_{\delta})}+\omega_{q}^{b_{1}(\vec{i}_{\delta})-b_{1}(\vec{i}^{\prime}_{\delta})}\right)\\ &\quad+\left(\omega_{q}^{a_{2}(\vec{i}_{\delta})-a_{1}(\vec{i}^{\prime}_{\delta})}+\omega_{q}^{b_{2}(\vec{i}_{\delta})-b_{1}(\vec{i}^{\prime}_{\delta})}\right)\\ &\quad+\left(\omega_{q}^{a_{2}(\vec{i}_{\delta})-a_{2}(\vec{i}^{\prime}_{\delta})}+\omega_{q}^{b_{2}(\vec{i}_{\delta})-b_{2}(\vec{i}^{\prime}_{\delta})}\right)\bigg{]}=0.\end{split}$ (82)
      using Lemma 7. Therefore
      $C\left(\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{1}},\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{2}}\right)(\tau_{1},\tau_{2})=0.$ (83)
    2. Sub-case 2:
      Let $\bm{\zeta}_{1}\neq\bm{\zeta}_{2}$. If $(\mathbf{s}_{i}^{1},\mathbf{t}_{i}^{1})=(\mathbf{s}_{i}^{2},\mathbf{t}_{i}^{2})$, then we proceed similar to the previous case and get
      $\begin{split}C\left(\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{1}},\mathbf{X}^{\bm{\gamma},\bm{\zeta}_{2}}\right)(\tau_{1},\tau_{2})=0.\end{split}$ (84)
      However, if either $\mathbf{s}_{i}^{1}\neq\mathbf{s}_{i}^{2}$ or $\mathbf{t}_{i}^{1}\neq\mathbf{t}_{i}^{2}$, then using Lemma 5, we have
      $\begin{split}\sum_{\bm{\gamma}\prod_{\alpha=1}^{k}\mathbb{Z}_{p_{\alpha}}}\sum_{j_{\xi}}\omega_{q}^{a_{\mathbf{s}_{i}^{1},\mathbf{t}_{i}^{1}}^{\bm{\gamma}}(\vec{j}_{\xi})-a_{\mathbf{s}_{i}^{2},\mathbf{t}_{i}^{2}}^{\bm{\gamma}}(\vec{j}_{\xi})}=0\end{split}.$ (85)
  3. Case III:
    Let $0\leq\tau_{1}<2^{m+1}$ and $-\prod_{\alpha=1}^{k}p_{\alpha}^{m_{\alpha}-1}<\tau_{2}<0$. This case can be proved similar to Case I.
Combining all these cases, we have the theorem.
Appendix C Proof of Theorem 2
[…]
$D^{\nu}_{i,j}=\begin{cases}A^{\nu}_{i,j}(t),&(t-1)(L_{1}+Z_{1}-1)\leq i<\\ &(t-1)(L_{1}+Z_{1}-1)+L_{1},\\ &0\leq j<L_{2},t\in\{1,2,\dots,M\};\\ 0,&\text{otherwise}.\end{cases}$ (86)
This produces $\sum_{i=0}^{N_{1}}\sum_{j=0}^{N_{2}}\left|D^{\nu}_{i,j}\right|^{2}=ML_{1}L_{2}$. Now, we apply cyclic shifts on the 2-D arrays $\mathbf{D}^{\nu}(\nu=1,2,\dots,K)$. We take $Z_{1}$ distinct cyclic shifts (including the 00 shift) in the right direction and $Z_{2}$ distinct cyclic shifts (including the 00 shift) in the downward direction and all their combinations. This procedure produces a new set $\mathcal{E}$ having $KZ_{1}Z_{2}$ distinct 2-D arrays of size $N_{1}\times N_{2}$. We apply Lemma 8 on this set =$\mathcal{E}=\{\mathbf{E}^{\nu}:\nu=1,2,\dots,KZ_{1}Z_{2}-1\}$. Then, we have
$\begin{split}&\overline{K}(\overline{K}\!\!-\!\!1)c^{2}_{max}+\overline{K}\left(ML_{1}L_{2}\right)^{2}\\ &\geq\!\!\!\sum_{s_{i}=0}^{\overline{L}_{1}-1}\sum_{s_{j}=0}^{\overline{L}_{2}-1}\left(\sum_{\nu=0}^{\overline{K}-1}\left|\mathbf{E}^{\nu}_{s_{i},s_{j}}\right|^{2}\right)^{2},\end{split}$ (87)
where $\overline{L}_{1}=M(L_{1}+Z_{1}-1)$, $\overline{L}_{2}=L_{2}+Z_{2}-1$ and $\overline{K}=KZ_{1}Z_{2}$. But, as we are considering a 2-D ZCACS, we would have $c_{max}=0$. Hence, we get
$\overline{K}\left(ML_{1}L_{2}\right)^{2}\geq\sum_{s_{i}=0}^{\overline{L}_{1}-1}\sum_{s_{j}=0}^{\overline{L}_{2}-1}\left(\sum_{\nu=0}^{\overline{K}-1}\left|\mathbf{E}^{\nu}_{s_{i},s_{j}}\right|^{2}\right)^{2}.$ (88)
Now, applying Cauchy-Schwarz inequality and changing the order of summation, we get
$\begin{split}&\sum_{s_{i}=0}^{\overline{L}_{1}-1}\sum_{s_{j}=0}^{\overline{L}_{2}-1}\left(\sum_{\nu=0}^{\overline{K}-1}\left|\mathbf{E}^{\nu}_{s_{i},s_{j}}\right|^{2}\right)^{2}\\ &\geq\frac{\left[\sum_{s_{i}=0}^{\overline{L}_{1}-1}\sum_{s_{j}=0}^{\overline{L}_{2}-1}\left(\sum_{\nu=0}^{\overline{K}-1}\left|\mathbf{E}^{\nu}_{s_{i},s_{j}}\right|^{2}\right)\right]^{2}}{\sum_{s_{i}=0}^{\overline{L}_{1}-1}\sum_{s_{j}=0}^{\overline{L}_{2}-1}1}\\ &=\frac{\left[\sum_{\nu=0}^{\overline{K}-1}ML_{1}L_{2}\right]^{2}}{ML_{1}L_{2}}=\overline{K}^{2}ML_{1}L_{2}.\end{split}$ (89)
From (88) and (89), we get the desired result.